the GRUB manual

The GRand Unified Bootloader, version 0.92, 29 April 2002.

Gordon Matzigkeit Yoshinori K. Okuji

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999,2000,2001,2002 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by Free Software Foundation.

1 Introduction to GRUB

1.1 Overview

Briefly, a boot loader is the first software program that runs when a computer starts. It is responsible for loading and transferring control to an operating system kernel software (such as Linux or GNU Mach). The kernel, in turn, initializes the rest of the operating system (e.g. a GNU system).

GNU GRUB is a very powerful boot loader, which can load a wide variety of free operating systems, as well as proprietary operating systems with chain-loading^{1}. GRUB is designed to address the complexity of booting a personal computer; both the program and this manual are tightly bound to that computer platform, although porting to other platforms may be addressed in the future.

One of the important features in GRUB is flexibility; GRUB understands filesystems and kernel executable formats, so you can load an arbitrary operating system the way you like, without recording the physical position of your kernel on the disk.

Thus you can load the kernel just by specifying its file name and the drive (and the partition) where the kernel resides. To let GRUB know the drive and the file name, you can either type in them manually via the command-line interface (see Section 12.1 [Commandline interface], page 20), or use the nice menu interface (see Section 12.2 [Menu interface], page 21) through which you can easily select which OS it boots. To allow you to customize the menu, GRUB will load a preexisting configuration file (see Chapter 5 [Configuration], page 11). Note that you can not only enter the command-line interface whenever you like, but also you can edit specific menu entries prior to using them.

In the following chapters, you will learn how to specify a drive or a partition, and a file name (see Chapter 2 [Naming convention], page 4) to GRUB, how to install GRUB on your drive (see Chapter 3 [Installation], page 5), and how to boot your OSes (see Chapter 4 [Booting], page 8), step by step.

1.2 History of GRUB

GRUB originated in 1995 when Erich Boleyn was trying to boot the GNU Hurd with the University of Utah’s Mach 4 microkernel (now known as GNU Mach). Erich and Brian Ford designed the Multiboot Specification (see section “Motivation” in The Multiboot Specification), because they were determined not to add to the large number of mutuallyincompatible PC boot methods.

Erich then began modifying the FreeBSD boot loader so that it would understand Multiboot. He soon realized that it would be a lot easier to write his own boot loader from scratch than to keep working on the FreeBSD boot loader, and so GRUB was born.

Erich added many features to GRUB, but other priorities prevented him from keeping up with the demands of its quickly-expanding user base. In 1999, Gordon Matzigkeit and Yoshinori K. Okuji adopted GRUB as an official GNU package, and opened its development by making the latest sources available via anonymous CVS. See Appendix B [Obtaining and Building GRUB], page 48, for more information.

1 chain-load is the mechanism for loading unsupported operating systems by loading another boot

loader. It is typically used for loading DOS or Windows.

1.3 GRUB features

The primary requirement for GRUB is that it be compliant with the Multiboot Specification, which is described in section “Motivation” in The Multiboot Specification.

The other goals, listed in approximate order of importance, are:

• Basic functions must be straightforward for end-users.

• Rich functionality to support kernel experts and designers.

• Backward compatibility for booting FreeBSD, NetBSD, OpenBSD, and Linux. Proprietary kernels (such as DOS, Windows NT, and OS/2) are supported via a chain-loading function.

Except for specific compatibility modes (chain-loading and the Linux piggyback format), all kernels will be started in much the same state as in the Multiboot Specification. Only kernels loaded at 1 megabyte or above are presently supported. Any attempt to load below that boundary will simply result in immediate failure and an error message reporting the problem.

In addition to the requirements above, GRUB has the following features (note that the Multiboot Specification doesn’t require all the features that GRUB supports):

Recognize multiple executable formats Support many of the a.out variants plus ELF. Symbol tables are also loaded.

Support non-Multiboot kernels Support many of the various free 32-bit kernels that lack Multiboot compliance (primarily FreeBSD, NetBSD, OpenBSD, and Linux). Chain-loading of other boot loaders is also supported.

Load multiples modules Fully support the Multiboot feature of loading multiple modules.

Load a configuration file Support a human-readable text configuration file with preset boot commands. You can also load another configuration file dynamically and embed a preset configuration file in a GRUB image file. The list of commands (see Chapter 13 [Commands], page 22) are a superset of those supported on the command-line. An example configuration file is provided in Chapter 5 [Configuration], page 11.

Provide a menu interface A menu interface listing the preset boot commands, with a programmable timeout, is available. There is no fixed limit on the number of boot entries, and the current implementation has space for several hundred.

Have a flexible command-line interface A fairly flexible command-line interface, accessible from the menu, is available to edit any preset commands, or write a new boot command set from scratch. If no configuration file is present, GRUB drops to the command-line.

The list of commands (see Chapter 13 [Commands], page 22) are a subset of those supported for configuration files. Editing commands closely resembles the Bash command-line (see section “Command Line Editing” in Bash Features), with ^{⟨}\underline{\overline{TAB}}^{⟩}-completion of commands, devices, partitions, and files in a directory depending on context.

Support multiple filesystem types Support multiple filesystem types transparently, plus a useful explicit blocklist notation. The currently supported filesystem types are BSD FFS, DOS FAT16 and FAT32, Minix fs, Linux ext2fs, ReiserFS, JFS, XFS, and VSTa fs. See Chapter 11 [Filesystem], page 19, for more information.

Support automatic decompression Can decompress files which were compressed by gzip. This function is both automatic and transparent to the user (i.e. all functions operate upon the uncompressed contents of the specified files). This greatly reduces a file size and the loading time, a particularly major benefit for floppies.^{2}

It is conceivable that some kernel modules should be loaded in a compressed state, so a different module-loading command can be specified to avoid uncompressing the modules.

Access data on any installed device Support reading data from any or all floppy or hard disk(s) recognized by the BIOS, independent of the setting of the root device.

Be independent of drive geometry translations Unlike many other boot loaders, GRUB makes the particular drive translation irrelevant. A drive installed and running with one translation may be converted to another translation without any adverse effects or changes in GRUB’s configuration.

Detect all installed ram GRUB can generally find all the installed ram on a PC-compatible machine. It uses an advanced BIOS query technique for finding all memory regions. As described on the Multiboot Specification (see section “Motivation” in The Multiboot Specification), not all kernels make use of this information, but GRUB provides it for those who do.

Support Logical Block Address mode In traditional disk calls (called CHS mode), there is a geometry translation problem, that is, the BIOS cannot access over 1024 cylinders, so the accessible space is limited to at least 508 MB and to at most 8GB. GRUB can’t universally solve this problem, as there is no standard interface used in all machines. However, several newer machines have the new interface, Logical Block Address (LBA) mode. GRUB automatically detects if LBA mode is available and uses it if available. In LBA mode, GRUB can access the entire disk.

2 There are a few pathological cases where loading a very badly organized ELF kernel might take

longer, but in practice this never happen.

Support network booting GRUB is basically a disk-based boot loader but also has network support. You can load OS images from a network by using the TFTP protocol.

Support remote terminals To support computers with no console, GRUB provides remote terminal support, so that you can control GRUB from a remote host. Only serial terminal support is implemented at the moment.

1.4 The role of a boot loader

The following is a quotation from Gordon Matzigkeit, a GRUB fanatic: Some people like to acknowledge both the operating system and kernel when they talk about their computers, so they might say they use “GNU/Linux” or “GNU/Hurd”. Other people seem to think that the kernel is the most important part of the system, so they like to call their GNU operating systems “Linux systems.” I, personally, believe that this is a grave injustice, because the boot loader is the most important software of all. I used to refer to the above systems as either “LILO”^{3} or “GRUB” systems. Unfortunately, nobody ever understood what I was talking about; now I just use the word “GNU” as a pseudonym for GRUB. So, if you ever hear people talking about their alleged “GNU” systems, remember that they are actually paying homage to the best boot loader around… GRUB! We, the GRUB maintainers, do not (usually) encourage Gordon’s level of fanaticism,

but it helps to remember that boot loaders deserve recognition. We hope that you enjoy using GNU GRUB as much as we did writing it.

2 Naming convention

The device syntax used in GRUB is a wee bit different from what you may have seen before in your operating system(s), and you need to know it so that you can specify a drive/partition.

Look at the following examples and explanations:

(fd0)

First of all, GRUB requires that the device name is enclosed with ‘(’ and ‘)’. The ‘fd’ part means that it is a floppy disk. The number ‘0’ is the drive number, which is counted from zero. This expression means that GRUB will use the whole floppy disk.

(hd0,1)

Here, ‘hd’ means it is a hard disk drive. The first integer ‘0’ indicates the drive number, that is, the first hard disk, while the second integer, ‘1’, indicates the partition number (or the pc slice number in the BSD terminology). Once again, please note that the partition numbers are counted from zero, not from one. This expression means the second partition of the first hard disk drive. In this case, GRUB uses one partition of the disk, instead of the whole disk.

3 The LInux LOader, a boot loader that everybody uses, but nobody likes.

     (hd0,4)

This specifies the first extended partition of the first hard disk drive. Note that the partition numbers for extended partitions are counted from ‘4’, regardless of the actual number of primary partitions on your hard disk.

     (hd1,a)

This means the BSD ‘a’ partition of the second hard disk. If you need to specify which pc slice number should be used, use something like this: ‘(hd1,0,a)’. If the pc slice number is omitted, GRUB searches for the first pc slice which has a BSD ‘a’ partition.

Of course, to actually access the disks or partitions with GRUB, you need to use the device specification in a command, like ‘root (fd0)’ or ‘unhide (hd0,2)’. To help you find out which number is a partition you want, the GRUB command-line (see Section 12.1 [Command-line interface], page 20) options have argument completion. That means that, for example, you only need to type ‘root (’, followed by a ^{⟨}\underline{\overline{TAB}}^{⟩}, and GRUB will display the list of drives, partitions, or file names, so it should be quite easy to determine the name of your target partition, even with minimal knowledge of the syntax.

Note that GRUB does not distinguish IDE from SCSI - it simply counts the drive numbers from zero, regardless of their type. Normally, any IDE drive number is less than any SCSI drive number, although that is not true if you change the boot sequence by swapping IDE and SCSI drives in your BIOS.

Now the question is, how to specify a file? Again, see this example:

     (hd0,0)/vmlinuz

This specifies the file named ‘vmlinuz’, found on the first partition of the first hard disk drive. Note that the argument completion works with file names, too.

That was easy, admit it. Do read the next chapter, to find out how to actually install GRUB on your drive.

3 Installation

First, you need to have GRUB itself properly installed on your system, (see Appendix B [Obtaining and Building GRUB], page 48) either from the source tarball, or as a package for your OS.

To use GRUB, you need to install it on your drive. There are two ways of doing that - either using the utility grub-install (see Chapter 16 [Invoking grub-install], page 44) on a UNIX-like OS, or by using the native Stage 2. These are quite similar, however, the utility might probe a wrong BIOS drive, so better be careful.

Also, if you install GRUB on a UNIX-like OS, please make sure that you have an emergency boot disk ready, so that you can rescue your computer if, by any chance, your hard drive becomes unusable (unbootable).

GRUB comes with boot images, which are normally installed in the directory

‘/usr/share/grub/i386-pc’. You need to copy the files ‘stage1’, ‘stage2’, and

‘*stage1_5’ to the directory ‘/boot/grub’. Here the directory where GRUB images are installed and the directory where GRUB will use to find them are called image directory and boot directory, respectively.

3.1 Creating a GRUB boot floppy

To create a GRUB boot floppy, you need to take the files ‘stage1’ and ‘stage2’ from the image directory, and write them to the first and the second block of the floppy disk, respectively.

Caution: This procedure will destroy any data currently stored on the floppy. On a UNIX-like operating system, that is done with the following commands:

     # cd /usr/share/grub/i386-pc
     # dd if=stage1 of=/dev/fd0 bs=512 count=1
     1+0 records in
     1+0 records out
     # dd if=stage2 of=/dev/fd0 bs=512 seek=1
     153+1 records in
     153+1 records out
     #

The device file name may be different. Consult the manual for your OS.

3.2 Installing GRUB natively

Caution: Installing GRUB’s stage1 in this manner will erase the normal boot-sector used by an OS.

GRUB can currently boot GNU Mach, Linux, FreeBSD, NetBSD, and OpenBSD directly, so using it on a boot sector should be okay. But generally, it would be a good idea to back up the first sector of the partition on which you are installing GRUB’s stage1. This isn’t as important if you are installing GRUB on the first sector of a hard disk, since it’s easy to reinitialize it (e.g. by running ‘FDISK /MBR’ from DOS).

If you decide to install GRUB in the native environment, which is definitely desirable, you’ll need to create the GRUB boot disk, and reboot your computer with it. Otherwise, see Section 3.3 [Installing GRUB using grub-install], page 7, for more details.

Once started, GRUB will show the command-line interface (see Section 12.1 [Command-line interface], page 20). First, set the GRUB’s root device^{4} to the boot directory, like this:

     grub> root (hd0,0)

If you are not sure which partition actually holds these files, use the command find (see Section 13.3.11 [find], page 31), like this:

     grub> find /boot/grub/stage1

This will search for the file name ‘/boot/grub/stage1’ and show the devices which contain the file.

4 Note that GRUB’s root device doesn’t necessarily mean your OS’s root partition; if you need to

specify a root partition for your OS, add the argument into the command kernel.

Once you’ve set the root device correctly, run the command setup (see Section 13.3.34 [setup], page 37):

     grub> setup (hd0)

This command will install GRUB on the MBR in the first drive. If you want to install GRUB into the boot sector of a partition instead of the MBR, specify a partition into which you want to install GRUB:

     grub> setup (hd0,0)

If you install GRUB into a partition or a drive other than the first one, you must chain-load GRUB from another boot loader. Refer to the manual for the boot loader to know how to chain-load GRUB.

Now you can boot GRUB without a GRUB floppy. See the chapter Chapter 4 [Booting], page 8 to find out how to boot your operating systems from GRUB.

3.3 Installing GRUB using grub-install

Caution: This procedure is definitely deprecated, because there are several posibilities that your computer can be unbootable. For example, most operating systems don’t tell GRUB how to map BIOS drives to OS devices correctly, GRUB merely guesses the mapping. This will succeed in most cases, but not always. So GRUB provides you with a user-defined map file called device map, which you must fix, if it is wrong. See Section 15.3 [Device map], page 43, for more details.

Unfortunately, if you do want to install GRUB under a UNIX-like OS (such as gnu), invoke the program grub-install (see Chapter 16 [Invoking grub-install], page 44) as the superuser (root).

The usage is basically very easy. You only need to specify one argument to the program, namely, where to install GRUB. The argument can be either of a device file or a GRUB’s drive/partition. So, this will install GRUB into the MBR of the first IDE disk under Linux:

     # grub-install /dev/hda

Likewise, under GNU/Hurd, this has the same effect:

     # grub-install /dev/hd0

If it is the first BIOS drive, this is the same as well:

     # grub-install '(hd0)'

But all the above examples assume that you use GRUB images under the root directory. If you want GRUB to use images under a directory other than the root directory, you need to specify the option ‘--root-directory’. The typical usage is that you create a GRUB boot floppy with a filesystem. Here is an example:

     # mke2fs /dev/fd0
     # mount -t ext2 /dev/fd0 /mnt
     # grub-install --root-directory=/mnt '(fd0)'
     # umount /mnt

Another example is in case that you have a separate boot partition which is mounted at ‘/boot’. Since GRUB is a boot loader, it doesn’t know anything about mountpoints at all. Thus, you need to run grub-install like this:

     # grub-install --root-directory=/boot /dev/hda

By the way, as noted above, it is quite difficult to guess BIOS drives correctly under a UNIX-like OS. Thus, grub-install will prompt you to check if it could really guess the correct mappings, after the installation. The format is defined in Section 15.3 [Device map], page 43. Please be careful enough. If the output is wrong, it is unlikely that your computer can boot with no problem.

Note that grub-install is actually just a shell script and the real task is done by the grub shell grub (see Chapter 15 [Invoking the grub shell], page 42). Therefore, you may run grub directly to install GRUB, without using grub-install. Don’t do that, however, unless you are very familiar with the internals of GRUB. Installing a boot loader on a running OS may be extremely dangerous.

4 Booting

For Multiboot-compliant kernels, GRUB can load them in a consistent way, but, for some free operating systems, you need to use some OS-specific magic.

4.1 How to boot operating systems

GRUB has two distinct boot methods. One of the two is to load an operating system directly, and the other is to chain-load another boot loader which then will load an operating system actually. Generally speaking, the former is desirable, because you don’t need to install or maintain other boot loaders and GRUB is flexible enough to load an operating system from an arbitrary disk/partition. However, the latter is sometimes required, since GRUB doesn’t support all the existing operating systems natively.

4.1.1 How to boot an OS directly with GRUB

Multiboot (see section “Motivation” in The Multiboot Specification) is the native format supported by GRUB. For the sake of convenience, there are also support for Linux, FreeBSD, NetBSD and OpenBSD. If you want to boot other operating systems, you will have to chain-load them (see Section 4.1.2 [Chain-loading], page 9).

Generally, GRUB can boot any Multiboot-compliant OS in the following steps:

1. Set GRUB’s root device to the drive where the OS images are stored by the command root (see Section 13.3.31 [root], page 36).

2. Load the kernel image by the command kernel (see Section 13.3.20 [kernel], page 33).

3. If you need modules, load them with the command module (see Section 13.3.25 [module], page 35) or modulenounzip (see Section 13.3.26 [modulenounzip], page 35).

4. Run the command boot (see Section 13.3.2 [boot], page 29).

Linux, FreeBSD, NetBSD and OpenBSD can be booted in a similar manner. You can load a kernel image by the command kernel and then run the command boot. If the kernel requires some parameters, just append the parameters to kernel, after the file name of the kernel. Also, please refer to Section 4.2 [OS-specific notes], page 9, for the information on your OS-specific issues.

4.1.2 Load another boot loader to boot unsupported operating systems

If you want to boot an unsupported operating system (e.g. Windows 95), chain-load a boot loader for the operating system. Normally, the boot loader is embedded in the boot sector of the partition on which the operating system is installed.

1. Set GRUB’s root device to the partition by the command rootnoverify (see Sec-

tion 13.3.32 [rootnoverify], page 36):

        grub> rootnoverify (hd0,0)

2. Set the active flag in the partition by the command makeactive^{5} (see Section 13.3.22

[makeactive], page 34):

        grub> makeactive

3. Load the boot loader by the command chainloader (see Section 13.3.4 [chainloader],

page 30):

        grub> chainloader +1

‘+1’ indicates that GRUB should read one sector from the start of the partition. The complete description about this syntax can be found in Section 11.3 [Block list syntax], page 20.

4. Run the command boot (see Section 13.3.2 [boot], page 29).

However, DOS and Windows have some deficiencies, so you might have to use more complicated instructions. See Section 4.2.6 [DOS/Windows], page 11, for more information.

4.2 Some caveats on OS-specific issues

Here, we describe some caveats on several operating systems.

4.2.1 GNU/Hurd

Since GNU/Hurd is Multiboot-compliant, it is easy to boot it; there is nothing special about it. But do not forget that you have to specify a root partition to the kernel.

1. Set GRUB’s root device to the same drive as GNU/Hurd’s. Probably the command find /boot/gnumach or similar can help you (see Section 13.3.11 [find], page 31).

2. Load the kernel and the module, like this:

        grub> kernel /boot/gnumach root=hd0s1
        grub> module /boot/serverboot

3. Run the command boot (see Section 13.3.2 [boot], page 29).

4.2.2 GNU/Linux

It is relatively easy to boot GNU/Linux from GRUB, because it somewhat resembles to boot a Multiboot-compliant OS.

1. Set GRUB’s root device to the same drive as GNU/Linux’s. Probably the command find /vmlinuz or similar can help you (see Section 13.3.11 [find], page 31).

5 This is not necessary for most of the modern operating systems.

2. Load the kernel:

         grub> kernel /vmlinuz root=/dev/hda1

If you need to specify some kernel parameters, just append them to the command. For example, to set ‘vga’ to ‘ext’, do this:

         grub> kernel /vmlinuz root=/dev/hda1 vga=ext

See the documentation in the Linux source tree for the complete information on the available options.

3. If you use an initrd, execute the command initrd (see Section 13.3.17 [initrd], page 32)

    after kernel:
         grub> initrd /initrd

4. Finally, run the command boot (see Section 13.3.2 [boot], page 29).

Caution: If you use an initrd and specify the ‘mem=’ option to the kernel, to let it use less than actual memory size, you will also have to specify the same memory size to GRUB. To let GRUB know the size, run the command uppermem before loading the kernel. See Section 13.3.37 [uppermem], page 37, for more information.

4.2.3 FreeBSD

GRUB can load the kernel directly, either in ELF or a.out format. But this is not recommended, since FreeBSD’s bootstrap interface sometimes changes heavily, so GRUB can’t guarantee to pass kernel parameters correctly.

Thus, we’d recommend loading the very flexible loader ‘/boot/loader’ instead. See this example:

     grub> root (hd0,a)
     grub> kernel /boot/loader
     grub> boot

4.2.4 NetBSD

GRUB can load NetBSD a.out and ELF directly, follow these steps:

1. Set GRUB’s root device with root (see Section 13.3.31 [root], page 36).

2. Load the kernel with kernel (see Section 13.3.20 [kernel], page 33). You should append

the ugly option ‘--type=netbsd’, if you want to load an ELF kernel, like this:

         grub> kernel --type=netbsd /netbsd-elf

3. Run boot (see Section 13.3.2 [boot], page 29).

For now, however, GRUB doesn’t allow you to pass kernel parameters, so it may be better to chain-load it instead, for more information please see Section 4.1.2 [Chain-loading], page 9.

4.2.5 OpenBSD

The booting instruction is exactly the same as for NetBSD (see Section 4.2.4 [NetBSD], page 10).

4.2.6 DOS/Windows

GRUB cannot boot DOS or Windows directly, so you must chain-load them (see Section 4.1.2 [Chain-loading], page 9). However, their boot loaders have some critical deficiencies, so it may not work to just chain-load them. To overcome the problems, GRUB provides you with two helper functions.

If you have installed DOS (or Windows) on a non-first hard disk, you have to use the disk swapping technique, because that OS cannot boot from any disks but the first one. The workaround used in GRUB is the command map (see Section 13.3.23 [map], page 34), like this:

     grub> map (hd0) (hd1)
     grub> map (hd1) (hd0)

This performs a virtual swap between your first and second hard drive.

Caution: This is effective only if DOS (or Windows) uses BIOS to access the swapped disks. If that OS uses a special driver for the disks, this probably won’t work.

Another problem arises if you installed more than one set of DOS/Windows onto one disk, because they could be confused if there are more than one primary partitions for DOS/Windows. Certainly you should avoid doing this, but there is a solution if you do want to do so. Use the partition hiding/unhiding technique.

If GRUB hides a DOS (or Windows) partition (see Section 13.2.5 [hide], page 25), DOS (or Windows) will ignore the partition. If GRUB unhides a DOS (or Windows) partition (see Section 13.2.16 [unhide], page 29), DOS (or Windows) will detect the partition. Thus, if you have installed DOS (or Windows) on the first and the second partition of the first hard disk, and you want to boot the copy on the first partition, do the following:

     grub> unhide (hd0,0)
     grub> hide (hd0,1)
     grub> rootnoverify (hd0,0)
     grub> chainloader +1
     grub> makeactive
     grub> boot

4.2.7 SCO UnixWare

It is known that the signature in the boot loader for SCO UnixWare is wrong, so you will have to specify the option ‘--force’ to chainloader (see Section 13.3.4 [chainloader], page 30), like this:

     grub> rootnoverify (hd1,0)
     grub> chainloader --force +1
     grub> makeactive
     grub> boot

5 Configuration

You probably noticed that you need to type several commands to boot your OS. There’s a solution to that - GRUB provides a menu interface (see Section 12.2 [Menu interface], page 21) from which you can select an item (using arrow keys) that will do everything to boot an OS.

To enable the menu, you need a configuration file, ‘menu.lst’ under the boot directory. We’ll analyze an example file.

The file first contains some general settings, the menu interface related options. You can put these commands (see Section 13.1 [Menu-specific commands], page 22) before any of the items (starting with title (see Section 13.1.5 [title], page 23)).

     #
     # Sample boot menu configuration file
     #

As you may have guessed, these lines are comments. Lines starting with a hash character (‘#’), and blank lines, are ignored by GRUB.

     # By default, boot the first entry.
     default 0

The first entry (here, counting starts with number zero, not one!) will be the default choice.

     # Boot automatically after 30 secs.
     timeout 30

As the comment says, GRUB will boot automatically in 30 seconds, unless interrupted with a keypress.

     # Fallback to the second entry.
     fallback 1

If, for any reason, the default entry doesn’t work, fall back to the second one (this is rarely used, for obvious reasons).

Note that the complete descriptions of these commands, which are menu interface specific, can be found in Section 13.1 [Menu-specific commands], page 22. Other descriptions can be found in Chapter 13 [Commands], page 22.

Now, on to the actual OS definitions. You will see that each entry begins with a special command, title (see Section 13.1.5 [title], page 23), and the action is described after it. Note that there is no command boot (see Section 13.3.2 [boot], page 29) at the end of each item. That is because GRUB automatically executes boot if it loads other commands successfully.

The argument for the command title is used to display a short title/description of the entry in the menu. Since title displays the argument as is, you can write basically anything in there.

     # For booting GNU/Hurd
     title  GNU/Hurd
     root   (hd0,0)
     kernel /boot/gnumach.gz root=hd0s1
     module /boot/serverboot.gz

This boots GNU/Hurd from the first hard disk.

     # For booting GNU/Linux
     title  GNU/Linux
     kernel (hd1,0)/vmlinuz root=/dev/hdb1

This boots GNU/Linux, but from the second hard disk.

# For booting Mach (getting kernel from floppy)
title  Utah Mach4 multiboot
root   (hd0,2)
pause  Insert the diskette now^G!!
kernel (fd0)/boot/kernel root=hd0s3
module (fd0)/boot/bootstrap

This boots Mach with a kernel on a floppy, but the root filesystem at hd0s3. It also contains a pause line (see Section 13.3.27 [pause], page 35), which will cause GRUB to display a prompt and delay, before actually executing the rest of the commands and booting.

# For booting FreeBSD
title  FreeBSD
root   (hd0,2,a)
kernel /boot/loader

This item will boot FreeBSD kernel loaded from the ‘a’ partition of the third pc slice of the first hard disk.

# For booting OS/2
title OS/2
root  (hd0,1)
makeactive
# chainload OS/2 bootloader from the first sector
chainloader +1
# This is similar to "chainload", but loads a specific file
#chainloader /boot/chain.os2

This will boot OS/2, using a chain-loader (see Section 4.1.2 [Chain-loading], page 9).

# For booting Windows NT or Windows95
title Windows NT / Windows 95 boot menu
root        (hd0,0)
makeactive
chainloader +1
# For loading DOS if Windows NT is installed
# chainload /bootsect.dos

The same as the above, but for Windows.

# For installing GRUB into the hard disk
title Install GRUB into the hard disk
root    (hd0,0)
setup   (hd0)

This will just (re)install GRUB onto the hard disk.

# Change the colors.
title Change the colors
color light-green/brown blink-red/blue

In the last entry, the command color is used (see Section 13.2.2 [color], page 24), to change the menu colors (try it!). This command is somewhat special, because it can be used both in the command-line and in the menu. GRUB has several such commands, see Section 13.2 [General commands], page 23.

We hope that you now understand how to use the basic features of GRUB. To learn more about GRUB, see the following chapters.

6 Downloading OS images from a network

Although GRUB is a disk-based boot loader, it does provide network support. To use the network support, you need to enable at least one network driver in the GRUB build process. For more information please see ‘netboot/README.netboot’ in the source distribution.

6.1 How to set up your network

GRUB requires a file server and optionally a server that will assign an IP address to the machine on which GRUB is running. For the former, only TFTP is supported at the moment. The latter is either BOOTP, DHCP or a RARP server^{6}. It is not necessary to run both the servers on one computer. How to configure these servers is beyond the scope of this document, so please refer to the manuals specific to those protocols/servers.

If you decided to use a server to assign an IP address, set up the server and run bootp (see Section 13.2.1 [bootp], page 23), dhcp (see Section 13.2.4 [dhcp], page 25) or rarp (see Section 13.2.11 [rarp], page 26) for BOOTP, DHCP or RARP, respectively. Each command will show an assigned IP address, a netmask, an IP address for your TFTP server and a gateway. If any of the addresses is wrong or it causes an error, probably the configuration of your servers isn’t set up properly.

Otherwise, run ifconfig, like this:

     grub> ifconfig --address=192.168.110.23 --server=192.168.110.14

You can also use ifconfig in conjugation with bootp, dhcp or rarp (e.g. to reassign the server address manually). See Section 13.2.6 [ifconfig], page 25, for more details.

Finally, download your OS images from your network. The network can be accessed using the network drive ‘(nd)’. Everything else is very similar to the normal instructions (see Chapter 4 [Booting], page 8).

Here is an example:

     grub> bootp
     Probing... [NE*000]
     NE2000 base ...
     Address: 192.168.110.23    Netmask: 255.255.255.0
     Server: 192.168.110.14     Gateway: 192.168.110.1
     grub> root (nd)
     grub> kernel /tftproot/gnumach.gz root=sd0s1
     grub> module /tftproot/serverboot.gz
     grub> boot

6.2 Booting from a network

It is sometimes very useful to boot from a network, especially, when you use a machine which has no local disk. In this case, you need to obtain a kind of Net Boot rom, such as a PXE rom or a free software package like Etherboot. Such a Boot rom first boots the machine, sets up the network card installed into the machine, and downloads a second stage boot image from the network. Then, the second image will try to boot an operating system from the network actually.

6 RARP is deprecated, since it cannot serve much information

GRUB provides two second stage images, ‘nbgrub’ and ‘pxegrub’ (see Chapter 10 [Images], page 18). Those images are the same as the normal Stage 2, except that they set up a network automatically, and try to load a configuration file from the network, if specified. The usage is very simple: If the machine has a PXE rom, use ‘pxegrub’. If the machine has a NBI loader such as Etherboot, use ‘nbgrub’. There is no difference between them but their formats. As how to load a second stage image you want to use should be described in the manual on your Net Boot rom, please refer to the manual, for more information.

However, there is one thing specific to GRUB. Namely, how to specify a configuration file in a BOOTP/DHCP server. For now, GRUB uses the tag ‘150’, to get the name of a configuration file. This below is an example about a BOOTP configuration:

     .allhost:hd=/tmp:bf=null:\
             :ds=145.71.35.1 145.71.32.1:\
             :sm=255.255.254.0:\
             :gw=145.71.35.1:\
             :sa=145.71.35.5:
     foo:ht=1:ha=63655d0334a7:ip=145.71.35.127:\
             :bf=/nbgrub:\
             :tc=.allhost:\
             :T150="(nd)/tftpboot/menu.lst.foo":

Note that you should specify the drive name (nd) in the name of the configuration file. That is because you can change the root drive before downloading the configuration from the TFTP server, when the preset menu feature is used (see Chapter 8 [Preset Menu], page 16).

See the manual about your BOOTP/DHCP server, for more information. The exact syntax should differ from the example, more or less.

7 Using GRUB via a serial line

This chapter describes how to use the serial terminal support in GRUB.

If you have many computers or computers with no display/keyboard, it would be very useful to control the computers with serial communications. To connect a computer with another via a serial line, you need to prepare a null-modem (cross) serial cable, and you may need to have multiport serial boards, if your computer doesn’t have extra serial ports. In addition, a terminal emulator is also required, such as minicom. Refer to a manual of your operating system, for more information.

As for GRUB, the instruction to set up a serial terminal is quite simple. First of all, make sure that you haven’t specified the option ‘--disable-serial’ to the configure script when you built your GRUB images. If you get them in binary form, probably they have serial terminal support already.

Then, initialize your serial terminal after GRUB starts up. Here is an example:

     grub> serial --unit=0 --speed=9600
     grub> terminal serial

The command serial initializes the serial unit 0 with the speed 9600bps. The serial unit 0 is usually called ‘COM1’, so, if you want to use COM2, you must specify ‘--unit=1’ instead. This command accepts many other options, so please refer to Section 13.2.12 [serial], page 26, for more details.

The command terminal (see Section 13.2.14 [terminal], page 28) chooses which type of terminal you want to use. In that case above, the terminal will be a serial terminal, but you can also pass console to the command, like ‘terminal serial console’. In this case, a terminal in which you press any key will be selected as a GRUB terminal.

However, note that GRUB assumes that your terminal emulator is compatible with VT100 by default. This is true for most terminal emulators nowadays, but you should pass the option ‘--dumb’ to the command, if your terminal emulator is not VT100-compatible or implements few VT100 escape sequences. If you specify the option, then GRUB provides you with an alternative menu interface, because the normal menu requires several fancy features for your terminal.

8 Embedding a configuration file into GRUB

GRUB supports preset menu which is always loaded before starting. The preset menu feature is useful, for example, when your computer has no console but a serial cable. In this case, it is critical to set up the serial terminal as soon as possible, since you cannot see any message until the serial terminal begins to work. So it is nice to run the commands serial (see Section 13.2.12 [serial], page 26) and terminal (see Section 13.2.14 [terminal], page 28) sooner than anything else at the start-up time.

It is slightly complicated how the preset menu works:

1. GRUB checks if the preset menu feature is used, and loads the preset menu, if available. This includes running commands and reading boot entries, like an ordinary configuration file.

2. GRUB checks if the configuration file is available. Note that this check is performed regardless of the existence of the preset menu. The configuration file is loaded, even after the preset menu was loaded.

3. When the preset menu includes any boot entries, they are cleared when the configuration file is loaded. It doesn’t matter whether the configuration file has any entries or no entry. The boot entries in the preset menu are used only when GRUB fails in loading the configuration file.

To enable the preset menu feature, you must specify a file to the configure script with the option ‘--enable-preset-menu’. The file has the same semantics as normal configuration files (see Chapter 5 [Configuration], page 11).

Another point you should take care is that the diskless support (see Section 6.2 [Diskless], page 14) diverts the preset menu. Diskless images embed a preset menu to execute the command bootp (see Section 13.2.1 [bootp], page 23) automatically, unless you specify your own preset menu to the configure script. This means that you must put commands to initialize a network in the preset menu yourself, because diskless images don’t set it up implicitly, when you use the preset menu explicitly.

Therefore, a typical preset menu used with diskless support would be like this:

     # Set up the serial terminal, first of all.
     serial --unit=0 --speed=19200
     terminal --timeout=0 serial
     # Initialize the network.
     dhcp

9 Protecting your computer from cracking

You may be interested in how to prevent ordinary users from doing whatever they like, if you share your computer with other people. So this chapter describes how to improve the security of GRUB.

One thing which could be a security hole is that the user can do too many things with GRUB, because GRUB allows to modify its configuration and run arbitrary commands at run-time. For example, the user can read even ‘/etc/passwd’ in the command-line interface by the command cat (see Section 13.3.3 [cat], page 30). So it is necessary to disable all the interactive operations.

Thus, GRUB provides password feature, so that only administrators can start the interactive operations (i.e. editing menu entries and entering the command-line interface). To use this feature, you need to run the command password in your configuration file (see Section 13.2.10 [password], page 26), like this:

     password --md5 PASSWORD

If this is specified, GRUB disallows any interactive control, until you press the key ⟨\underline{\overline{p}}^{⟩} and enter a correct password. The option ‘--md5’ tells GRUB that ‘PASSWORD’ is in MD5 format. If it is omitted, GRUB assumes the ‘PASSWORD’ is in clear text.

You can encrypt your password with the command md5crypt (see Section 13.3.24 [md5crypt], page 35). For example, run the grub shell (see Chapter 15 [Invoking the grub shell], page 42), and enter your password:

     grub> md5crypt
     Password: **********
     Encrypted: $1$U$JK7xFegdxWH6VuppCUSIb.

Then, cut and paste the encrypted password to your configuration file. Also, you can specify an optional argument to password. See this example:

     password PASSWORD /boot/grub/menu-admin.lst

In this case, GRUB will load ‘/boot/grub/menu-admin.lst’ as a configuration file when you enter the valid password.

Another thing which may be dangerous is that any user can choose any menu entry. Usually, this wouldn’t be problematic, but you might want to permit only administrators to run some of your menu entries, such as an entry for booting an insecure OS like DOS.

GRUB provides the command lock (see Section 13.3.21 [lock], page 34). This command always fails until you enter a valid password, so you can use it, like this:

     title Boot DOS
     lock
     rootnoverify (hd0,1)
     makeactive
     chainload +1

You should insert lock right after title, because any user can execute commands in an entry, until GRUB encounters lock.

You can also use the command password instead of lock. In this case the boot process will ask for the password and stop if it was entered incorrectly. Since the password takes its own PASSWORD argument this is useful if you want different passwords for different entries.

10 GRUB image files

GRUB consists of several images: two essential stages, optional stages called Stage 1.5, and two network boot images. Here is a short overview of them. See Chapter 20 [Internals], page 50, for more details.

‘stage1’ This is an essential image used for booting up GRUB. Usually, this is embedded in a MBR or the boot sector of a partition. Because a PC boot sector is 512 bytes, the size of this image is exactly 512 bytes. All ‘stage1’ must do is to load Stage 2 or Stage 1.5 from a local disk. Because of the size restriction, ‘stage1’ encodes the location of Stage 2 (or Stage 1.5) in a block list format, so it never understand any filesystem structure.

‘stage2’ This is the core image of GRUB. This does all things but booting up itself. Usually, this is put in a filesystem, but that is not required.

‘e2fs_stage1_5’
‘fat_stage1_5’
‘ffs_stage1_5’
‘jfs_stage1_5’
‘minix_stage1_5’
‘reiserfs_stage1_5’
‘vstafs_stage1_5’
‘xfs_stage1_5’

These are called Stage 1.5, because the purpose is a bridge between ‘stage1’ and ‘stage2’, that is to say, Stage 1.5 is loaded by Stage 1 and Stage 1.5 loads Stage 2. The difference between ‘stage1’ and ‘*_stage1_5’ is that the former doesn’t understand any filesystem but the latter does an filesystem (e.g. ‘e2fs_stage1_5’ understands ext2fs). So you can move the location of Stage 2 to another safely, even after GRUB has been installed. While Stage 2 cannot generally be embedded in a fixed area as the size is so large, Stage 1.5 can be installed into the area right after a MBR, or the boot loader area of a ReiserFS or a FFS.

‘nbgrub’ This is a network boot image for the Network Image Proposal used by some network boot loaders, such as Etherboot. This is mostly the same as Stage 2, but this also sets up a network and loads a configuration file from the network.

‘pxegrub’ This is another network boot image for the Preboot Execution Environment used by several Netboot ROMs. This is identical to ‘nbgrub’, except for the format.

11 Filesystem syntax and semantics

GRUB uses a special syntax for specifying disk drives which can be accessed by BIOS. Because of BIOS limitations, GRUB cannot distinguish between IDE, ESDI, SCSI, or others. You must know yourself which BIOS device is equivalent to which OS device. Normally, that will be clear if you see the files in a device or use the command find (see Section 13.3.11 [find], page 31).

11.1 How to specify devices

The device syntax is like this:

(device[,part-num][,bsd-subpart-letter])

‘[]’ means the parameter is optional. device should be either ‘fd’ or ‘hd’ followed by a digit, like ‘fd0’. But you can also set device to a hexadecimal or a decimal, which is a BIOS drive number, so the following are equivalent:

     (hd0)
     (0x80)
     (128)

part-num represents the partition number of device, starting from zero for primary partitions and from four for extended partitions, and bsd-subpart-letter represents the BSD disklabel subpartition, such as ‘a’ or ‘e’.

A shortcut for specifying BSD subpartitions is (device,bsd-subpart-letter), in this case, GRUB searches for the first PC partition containing a BSD disklabel, then finds the subpartition bsd-subpart-letter. Here is an example:

     (hd0,a)

The syntax like ‘(hd0)’ represents using the entire disk (or the MBR when installing GRUB), while the syntax like ‘(hd0,0)’ represents using the partition of the disk (or the boot sector of the partition when installing GRUB).

If you enabled the network support, the special drive, ‘(nd)’, is also available. Before using the network drive, you must initialize the network. See Chapter 6 [Network], page 14, for more information.

11.2 How to specify files

There are two ways to specify files, by absolute file name and by block list.

An absolute file name resembles a Unix absolute file name, using ‘/’ for the directory separator (not ‘\’ as in DOS). One example is ‘(hd0,0)/boot/grub/menu.lst’. This means the file ‘/boot/grub/menu.lst’ in the first partition of the first hard disk. If you omit the device name in an absolute file name, GRUB uses GRUB’s root device implicitly. So if you set the root device to, say, ‘(hd1,0)’ by the command root (see Section 13.3.31 [root],

page 36), then /boot/kernel is the same as (hd1,0)/boot/kernel.

11.3 How to specify block lists

A block list is used for specifying a file that doesn’t appear in the filesystem, like a chainloader. The syntax is [offset]+length[,[offset]+length].... Here is an example:

     0+100,200+1,300+300

This represents that GRUB should read blocks 0 through 99, block 200, and blocks 300 through 599. If you omit an offset, then GRUB assumes the offset is zero.

Like the file name syntax (see Section 11.2 [File name syntax], page 19), if a blocklist does not contain a device name, then GRUB uses GRUB’s root device. So (hd0,1)+1 is the same as +1 when the root device is ‘(hd0,1)’.

12 GRUB’s user interface

GRUB has both a simple menu interface for choosing preset entries from a configuration file, and a highly flexible command-line for performing any desired combination of boot commands.

GRUB looks for its configuration file as soon as it is loaded. If one is found, then the full menu interface is activated using whatever entries were found in the file. If you choose the command-line menu option, or if the configuration file was not found, then GRUB drops to the command-line interface.

12.1 The flexible command-line interface

The command-line interface provides a prompt and after it an editable text area much like a command-line in Unix or DOS. Each command is immediately executed after it is entered^{7}. The commands (see Section 13.3 [Command-line and menu entry commands], page 29) are a subset of those available in the configuration file, used with exactly the same syntax.

Cursor movement and editing of the text on the line can be done via a subset of the functions available in the Bash shell:

⟨\underline{\overline{C-f}}⟩
⟨\underline{\overline{PC right key}}⟩

Move forward one character.

⟨\underline{\overline{C-b}}⟩

⟨\underline{\overline{PC left key}}^{⟩} Move back one character.

⟨\underline{\overline{C-a}}⟩

⟨\underline{\overline{HOME}}^{⟩} Move to the start of the line.

⟨\underline{\overline{C-e}}⟩

⟨\underline{\overline{END}}^{⟩} Move the the end of the line.

⟨\underline{\overline{C-d}}⟩

⟨\underline{\overline{DEL}}^{⟩} Delete the character underneath the cursor.

7 However, this behavior will be changed in the future version, in a user-invisible way.

⟨\underline{\overline{C-h}}⟩

⟨\underline{\overline{BS}}^{⟩} Delete the character to the left of the cursor.

⟨\underline{\overline{C-k}}^{⟩} Kill the text from the current cursor position to the end of the line.

⟨\underline{\overline{C-u}}^{⟩} Kill backward from the cursor to the beginning of the line.

⟨\underline{\overline{C-y}}^{⟩} Yank the killed text back into the buffer at the cursor.

⟨\underline{\overline{C-p}}⟩

⟨\underline{\overline{PC up key}}^{⟩} Move up through the history list.

⟨\underline{\overline{C-n}}⟩

⟨

PC down key

⟩

Move down through the history list.

When typing commands interactively, if the cursor is within or before the first word in the command-line, pressing the ^{⟨}\underline{\overline{TAB}}^{⟩} key (or ^{⟨}\underline{\overline{C-i}}^{⟩}) will display a listing of the available commands, and if the cursor is after the first word, the ^{⟨}\underline{\overline{TAB}}^{⟩} will provide a completion listing of disks, partitions, and file names depending on the context.

Note that you cannot use the completion functionality in the TFTP filesystem. This is because TFTP doesn’t support file name listing for the security.

12.2 The simple menu interface

The menu interface is quite easy to use. Its commands are both reasonably intuitive and described on screen.

Basically, the menu interface provides a list of boot entries to the user to choose from. Use the arrow keys to select the entry of choice, then press ^{⟨}\underline{\overline{RET}}^{⟩} to run it. An optional timeout is available to boot the default entry (the first one if not set), which is aborted by pressing any key.

Commands are available to enter a bare command-line by pressing^{⟨}\underline{\overline{c}}^{⟩} (which operates exactly like the non-config-file version of GRUB, but allows one to return to the menu if desired by pressing ^{⟨}\underline{\overline{ESC}}^{⟩}) or to edit any of the boot entries by pressing ^{⟨}\underline{\overline{e}}^{⟩}.

If you protect the menu interface with a password (see Chapter 9 [Security], page 17), all you can do is choose an entry by pressing ^{⟨}\underline{\overline{RET}}^{⟩}, or press ^{⟨}\underline{\overline{p}}^{⟩} to enter the password.

12.3 Editing a menu entry

The menu entry editor looks much like the main menu interface, but the lines in the menu are individual commands in the selected entry instead of entry names.

If an ^{⟨}\underline{\overline{ESC}}^{⟩} is pressed in the editor, it aborts all the changes made to the configuration entry and returns to the main menu interface.

When a particular line is selected, the editor places the user at a special version of the GRUB command-line to edit that line. When the user hits ^{⟨}\underline{\overline{RET}}^{⟩}, GRUB replaces the line in question in the boot entry with the changes (unless it was aborted via ^{⟨}\underline{\overline{ESC}}^{⟩}, in which case the changes are thrown away).

If you want to add a new line to the menu entry, press ^{⟨}\underline{\overline{o}}^{⟩} if adding a line after the current line or press ^{⟨}\underline{\overline{O}}^{⟩} if before the current line.

To delete a line, hit the key ^{⟨}\underline{\overline{d}}^{⟩}. Although GRUB does not support undo unfortunately, you can do almost the same thing by just returning to the main menu.

12.4 The hidden menu interface

When your terminal is dumb or you request GRUB of hiding the menu interface explicitly with the command hiddenmenu (see Section 13.1.3 [hiddenmenu], page 23), GRUB doesn’t show the menu interface (see Section 12.2 [Menu interface], page 21) and automatically boots the default entry, unless interrupted by pressing ^{⟨}\underline{\overline{ESC}}^{⟩}.

When you interrupt the timeout and your terminal is dumb, GRUB falls back to the command-line interface (see Section 12.1 [Command-line interface], page 20).

13 The list of available commands

In this chapter, we list all commands that are available in GRUB.

Commands belong to different groups. A few can only be used in the global section of the configuration file (or “menu”); most of them can be entered on the command-line and can be either used in the menu or in the menu entries.

13.1 The list of commands for the menu only

The semantics used in parsing the configuration file are the following:

• The menu-specific commands have to be used before any others.

• The files must be in plain-text format.

• ‘#’ at the beginning of a line in a configuration file means it is only a comment.

• Options are separated by spaces.

• All numbers can be either decimal or hexadecimal. A hexadecimal number must be preceded by ‘0x’, and is case-insensitive.

• Extra options or text at the end of the line is ignored unless otherwise specified.

• Unrecognized commands are added to the current entry, except before entries start, where they are ignored.

These commands can only be used in the menu:

13.1.1 default

Command default num

Set the default entry to the entry number num. Numbering starts from 0, and the entry number 0 is the default if the command is not used. You can specify ‘saved’ instead of a number. In this case, the default entry is the entry saved with the command savedefault. See Section 13.3.33 [savedefault], page 36, for more information.

13.1.2 fallback

Command fallback num

Go into unattended boot mode: if the default boot entry has any errors, instead of waiting for the user to do anything, immediately start over using the num entry (same numbering as the default command (see Section 13.1.1 [default], page 22)). This obviously won’t help if the machine was rebooted by a kernel that GRUB loaded.

13.1.3 hiddenmenu

Command hiddenmenu

Don’t display the menu. If the command is used, no menu will be displayed on the control terminal, and the default entry will be booted after the timeout expired. The user can still request the menu to be displayed by pressing ^{⟨}\underline{\overline{ESC}}^{⟩} before the timeout expires. See also Section 12.4 [Hidden menu interface], page 22.

13.1.4 timeout

Command timeout sec

Set a timeout, in sec seconds, before automatically booting the default entry (normally the first entry defined).

13.1.5 title

Command title name …

Start a new boot entry, and set its name to the contents of the rest of the line, starting with the first non-space character.

13.2 The list of general commands

Commands usable both in the menu and in the command-line.

13.2.1 bootp

Command bootp [‘--with-configfile’]

Initialize a network device via the BOOTP protocol. This command is only available if GRUB is compiled with netboot support. See also Chapter 6 [Network], page 14.

If you specify ‘--with-configfile’ to this command, GRUB will fetch and load a configuration file specified by your BOOTP server with the vendor tag ‘150’.

13.2.2 color

Command color normal [highlight] Change the menu colors. The color normal is used for most lines in the menu (see Section 12.2 [Menu interface], page 21), and the color highlight is used to highlight the line where the cursor points. If you omit highlight, then the inverted color of normal is used for the highlighted line. The format of a color is foreground/background. foreground and background are symbolic color names. A symbolic color name must be one of these:

• black

• blue

• green

• cyan

• red

• magenta

• brown

• light-gray These below can be specified only for the foreground.

• dark-gray

• light-blue

• light-green

• light-cyan

• light-red

• light-magenta

• yellow

• white But only the first eight names can be used for background. You can prefix blink- to foreground if you want a blinking foreground color. This command can be used in the configuration file and on the command line, so you may write something like this in your configuration file:

   # Set default colors.
   color light-gray/blue black/light-gray
   # Change the colors.
   title OS-BS like
   color magenta/blue black/magenta

13.2.3 device

Command device drive file In the grub shell, specify the file file as the actual drive for a bios drive drive. You can use this command to create a disk image, and/or to fix the drives guessed by GRUB when GRUB fails to determine them correctly, like this:

     grub> device (fd0) /floppy-image
     grub> device (hd0) /dev/sd0

This command can be used only in the grub shell (see Chapter 15 [Invoking the grub shell], page 42).

13.2.4 dhcp

Command dhcp [–with-configfile] Initialize a network device via the DHCP protocol. Currently, this command is just an alias for bootp, since the two protocols are very similar. This command is only available if GRUB is compiled with netboot support. See also Chapter 6 [Network], page 14. If you specify ‘--with-configfile’ to this command, GRUB will fetch and load a configuration file specified by your DHCP server with the vendor tag ‘150’.

13.2.5 hide

Command hide partition Hide the partition partition by setting the hidden bit in its partition type code. This is useful only when booting DOS or Windows and multiple primary FAT partitions exist in one disk. See also Section 4.2.6 [DOS/Windows], page 11.

13.2.6 ifconfig

Command ifconfig [‘--server=server’] [‘--gateway=gateway’] [‘--mask=mask’]
    [‘--address=address’]

Configure the IP address, the netmask, the gateway, and the server address of a network device manually. The values must be in dotted decimal format, like ‘192.168.11.178’. The order of the options is not important. This command shows current network configuration, if no option is specified. See also Chapter 6 [Network], page 14.

13.2.7 pager

Command pager [flag]

Toggle or set the state of the internal pager. If flag is ‘on’, the internal pager is enabled. If flag is ‘off’, it is disabled. If no argument is given, the state is toggled.

13.2.8 partnew

Command partnew part type from to Create a new primary partition. part is a partition specification in GRUB syntax (see Chapter 2 [Naming convention], page 4); type is the partition type and must be a number in the range 0-0xff; from and to are the starting and ending sectors, expressed as an absolute sector number.

13.2.9 parttype

Command parttype part type

Change the type of an existing partition. part is a partition specification in GRUB syntax (see Chapter 2 [Naming convention], page 4); type is the new partition type and must be a number in the range 0-0xff.

13.2.10 password

Command password [‘--md5’] passwd [new-config-file] If used in the first section of a menu file, disable all interactive editing control (menu entry editor and command-line) and entries protected by the command lock. If the password passwd is entered, it loads the new-config-file as a new config file and restarts the GRUB Stage 2, if new-config-file is specified. Otherwise, GRUB will just unlock the privileged instructions. You can also use this command in the script section, in which case it will ask for the password, before continueing. The option ‘--md5’ tells GRUB that passwd is encrypted with md5crypt (see Section 13.3.24 [md5crypt], page 35).

13.2.11 rarp

Command rarp

Initialize a network device via the RARP protocol. This command is only available if GRUB is compiled with netboot support. See also Chapter 6 [Network], page 14.

13.2.12 serial

Command serial [‘--unit=unit’] [‘--port=port’] [‘--speed=speed’]
    [‘--word=word’] [‘--parity=parity’] [‘--stop=stop’] [‘--device=dev’]

Initialize a serial device. unit is a number in the range 0-3 specifying which serial port to use; default is 0, that corresponds the port often called COM1. port is the I/O port where the UART is to be found; if specified it takes precedence over unit. speed is the transmission speed; default is 9600. word and stop are the number of data bits and stop bits. Data bits must be in the range 5-8 and stop bits are 1 or 2. Default is 8 data bits and one stop bit. parity is one of ‘no’, ‘odd’, ‘even’ and defaults to ‘no’. The option ‘--device’ can only be used in the grub shell and is used to specify the tty device to be used in the host operating system (see Chapter 15 [Invoking the grub shell], page 42). The serial port is not used as a communication channel unless the terminal command is used (see Section 13.2.14 [terminal], page 28). This command is only available if GRUB is compiled with serial support. See also Chapter 7 [Serial terminal], page 15.

13.2.13 setkey

Command setkey [to key from key] Change the keyboard map. The key from key is mapped to the key to key. If no argument is specified, reset key mappings. Note that this command does not exchange the keys. If you want to exchange the keys, run this command again with the arguments exchanged, like this:

     grub> setkey capslock control
     grub> setkey control capslock

A key must be an alphabet, a digit, or one of these symbols: ‘escape’, ‘exclam’, ‘at’,

‘numbersign’, ‘dollar’, ‘percent’, ‘caret’, ‘ampersand’, ‘asterisk’, ‘parenleft’,
‘parenright’, ‘minus’, ‘underscore’, ‘equal’, ‘plus’, ‘backspace’, ‘tab’,
‘bracketleft’, ‘braceleft’, ‘bracketright’, ‘braceright’, ‘enter’, ‘control’,
‘semicolon’, ‘colon’, ‘quote’, ‘doublequote’, ‘backquote’, ‘tilde’, ‘shift’,
‘backslash’, ‘bar’, ‘comma’, ‘less’, ‘period’, ‘greater’, ‘slash’, ‘question’, ‘alt’,

‘space’, ‘capslock’, ‘FX’ (‘X’ is a digit), and ‘delete’. This table describes to which character each of the symbols corresponds:

‘exclam’ ‘!’

‘at’ ‘@’

‘numbersign’

‘#’

‘dollar’ ‘$’
‘percent’ ‘%’
‘caret’  ‘^’
‘ampersand’

‘&’

‘asterisk’

‘*’

‘parenleft’

‘(’

‘parenright’

‘)’

‘minus’  ‘-’
‘underscore’

‘_’

‘equal’  ‘=’
‘plus’  ‘+’
‘bracketleft’

‘[’

‘braceleft’

‘{’

‘bracketright’

‘]’

‘braceright’

‘}’

‘semicolon’

‘;’

‘colon’  ‘:’
‘quote’  ‘'’
‘doublequote’

‘"’

‘backquote’

‘`’

‘tilde’  ‘~’
‘backslash’

‘\’

‘bar’  ‘|’
‘comma’  ‘,’
‘less’  ‘<’
‘period’ ‘.’
‘greater’ ‘>’
‘slash’  ‘/’
‘question’

‘?’

‘space’  ‘ ’

13.2.14 terminal

Command terminal [‘--dumb’] [‘--timeout=secs’] [‘--lines=lines’] [‘--silent’]
    [‘console’] [‘serial’] [‘hercules’]

Select a terminal for user interaction. The terminal is assumed to be VT100compatible unless ‘--dumb’ is specified. If both ‘console’ and ‘serial’ are specified, then GRUB will use the one where a key is entered first or the first when the timeout expires. If neither are specified, the current setting is reported. This command is only available if GRUB is compiled with serial support. See also Chapter 7 [Serial terminal], page 15.

This may not make sense for most users, but GRUB supports Hercules console as well. Hercules console is usable like the ordinary console, and the usage is quite similar to that for serial terminals: specify ‘hercules’ as the argument. The option ‘--lines’ defines the number of lines in your terminal, and it is used for the internal pager function. If you don’t specify this option, the number is assumed as 24. The option ‘--silent’ suppresses the message to prompt you to hit any key. This might be useful if your system has no terminal device.

13.2.15 tftpserver

Command tftpserver ipaddr

Caution: This command exists only for backward compatibility. Use ifconfig (see Section 13.2.6 [ifconfig], page 25) instead. Override a TFTP server address returned by a BOOTP/DHCP/RARP server. The argument ipaddr must be in dotted decimal format, like ‘192.168.0.15’. This command is only available if GRUB is compiled with netboot support. See also Chapter 6 [Network], page 14.

13.2.16 unhide

Command unhide partition Unhide the partition partition by clearing the hidden bit in its partition type code. This is useful only when booting DOS or Windows and multiple primary partitions exist in one disk. See also Section 4.2.6 [DOS/Windows], page 11.

13.3 The list of command-line and menu entry commands

These commands are usable in the command-line and in menu entries. If you forget a command, you can run the command help (see Section 13.3.15 [help], page 32).

13.3.1 blocklist

Command blocklist file

Print the block list notation of the file file. See Section 11.3 [Block list syntax], page 20.

13.3.2 boot

Command boot

Boot the OS/chain-loader which has been loaded. Only necessary if running the fully interactive command-line (it is implicit at the end of a menu entry).

13.3.3 cat

Command cat file

Display the contents of the file file. This command may be useful to remind you of your OS’s root partition:

     grub> cat /etc/fstab

13.3.4 chainloader

Command chainloader [‘--force’] file Load file as a chain-loader. Like any other file loaded by the filesystem code, it can use the blocklist notation to grab the first sector of the current partition with ‘+1’. If you specify the option ‘--force’, then load file forcibly, whether it has a correct signature or not. This is required when you want to load a defective boot loader, such as SCO UnixWare 7.1 (see Section 4.2.7 [SCO UnixWare], page 11).

13.3.5 cmp

Command cmp file1 file2

Compare the file file1 with the file file2. If they differ in size, print the sizes like this:

     Differ in size: 0x1234 [foo], 0x4321 [bar]

If the sizes are equal but the bytes at an offset differ, then print the bytes like this:

     Differ at the offset 777: 0xbe [foo], 0xef [bar]

If they are completely identical, nothing will be printed.

13.3.6 configfile

Command configfile file

Load file as a configuration file.

13.3.7 debug

Command debug

Toggle debug mode (by default it is off). When debug mode is on, some extra messages are printed to show disk activity. This global debug flag is mainly useful for GRUB developers when testing new code.

13.3.8 displayapm

Command displayapm

Display APM BIOS information.

13.3.9 displaymem

Command displaymem

Display what GRUB thinks the system address space map of the machine is, including all regions of physical ram installed. GRUB’s upper/lower memory display uses the standard BIOS interface for the available memory in the first megabyte, or lower memory, and a synthesized number from various BIOS interfaces of the memory starting at 1MB and going up to the first chipset hole for upper memory (the standard PC upper memory interface is limited to reporting a maximum of 64MB).

13.3.10 embed

Command embed stage1 5 device Embed the Stage 1.5 stage1 5 in the sectors after the MBR if device is a drive, or in the boot loader area if device is a FFS partition or a ReiserFS partition.^{8} Print the number of sectors which stage1 5 occupies, if successful.

Usually, you don’t need to run this command directly. See Section 13.3.34 [setup], page 37.

13.3.11 find

Command find filename Search for the file name filename in all of partitions and print the list of the devices which contain the file. The file name filename should be an absolute file name like

/boot/grub/stage1.

13.3.12 fstest

Command fstest

Toggle filesystem test mode. Filesystem test mode, when turned on, prints out data corresponding to all the device reads and what values are being sent to the low-level routines. The format is ‘<partition-offset-sector, byte-offset, byte-length>’ for highlevel reads inside a partition, and ‘[disk-offset-sector]’ for low-level sector requests from the disk. Filesystem test mode is turned off by any use of the install (see Section 13.3.18 [install], page 32) or testload (see Section 13.3.35 [testload], page 37) commands.

13.3.13 geometry

8 The latter feature has not been implemented yet.

Command geometry drive [cylinder head sector [total sector]] Print the information for the drive drive. In the grub shell, you can set the geometry of the drive arbitrarily. The number of the cylinders, the one of the heads, the one of the sectors and the one of the total sectors are set to CYLINDER, HEAD, SECTOR and TOTAL SECTOR, respectively. If you omit TOTAL SECTOR, then it will be calculated based on the C/H/S values automatically.

13.3.14 halt

Command halt ‘--no-apm’ The command halts the computer. If the ‘--no-apm’ option is specified, no APM BIOS call is performed. Otherwise, the computer is shut down using APM.

13.3.15 help

Command help ‘--all’ [pattern …] Display helpful information about builtin commands. If you do not specify pattern, this command shows short descriptions of most of available commands. If you specify the option ‘--all’ to this command, short descriptions of rarely used commands (such as Section 13.3.35 [testload], page 37) are displayed as well. If you specify any patterns, it displays longer information about each of the commands which match those patterns.

13.3.16 impsprobe

Command impsprobe

Probe the Intel Multiprocessor Specification 1.1 or 1.4 configuration table and boot the various CPUs which are found into a tight loop. This command can be used only in the Stage 2.

13.3.17 initrd

Command initrd file …

Load an initial ramdisk for a Linux format boot image and set the appropriate parameters in the Linux setup area in memory. See also Section 4.2.2 [GNU/Linux], page 9.

13.3.18 install

Command install [‘--force-lba’] [‘--stage2=os_stage2_file’] stage1 file [‘d’]

dest dev stage2 file [addr] [‘p’] [config file] [real config file] This command is fairly complex, and you should not use this command unless you are familiar with GRUB. Use setup (see Section 13.3.34 [setup], page 37) instead.

In short, it will perform a full install presuming the Stage 2 or Stage 1.5^{9} is in its final install location. In slightly more detail, it will load stage1 file, validate that it is a GRUB Stage 1 of the right version number, install a blocklist for loading stage2 file as a Stage 2. If the option ‘d’ is present, the Stage 1 will always look for the actual disk stage2 file was installed on, rather than using the booting drive. The Stage 2 will be loaded at address addr, which must be ‘0x8000’ for a true Stage 2, and ‘0x2000’ for a Stage 1.5. If addr is not present, GRUB will determine the address automatically. It then writes the completed Stage 1 to the first block of the device dest dev. If the options ‘p’ or config file are present, then it reads the first block of stage2, modifies it with the values of the partition stage2 file was found on (for ‘p’) or places the string config file into the area telling the stage2 where to look for a configuration file at boot time. Likewise, if real config file is present and stage2 file is a Stage 1.5, then the Stage 2 config file is patched with the configuration file name real config file. This command preserves the DOS BPB (and for hard disks, the partition table) of the sector the Stage 1 is to be installed into. Caution: Several buggy BIOSes don’t pass a booting drive properly when booting from a hard disk drive. Therefore, you will have to specify the option ‘d’, whether your Stage2 resides at the booting drive or not, if you have such a BIOS unfortunately. We know these are defective in that:

Fujitsu LifeBook 400 BIOS version 31J0103A

HP Vectra XU 6/200 BIOS version GG.06.11

Caution2: A number of BIOSes don’t return a correct LBA support bitmap even if they do have the support. So GRUB provides a solution to ignore the wrong bitmap, that is, the option ‘--force-lba’. Don’t use this option if you know that your BIOS doesn’t have LBA support. Caution3: You must specify the option ‘--stage2’ in the grub shell, if you cannot unmount the filesystem where your stage2 file resides. The argument should be the file name in your operating system.

13.3.19 ioprobe

Command ioprobe drive

Probe I/O ports used for the drive drive. This command will list the I/O ports on the screen. For technical information, See Chapter 20 [Internals], page 50.

13.3.20 kernel

Command kernel [‘--type=type’] [‘--no-mem-option’] file … Attempt to load the primary boot image (Multiboot a.out or elf, Linux zImage or bzImage, FreeBSD a.out, NetBSD a.out, etc.) from file. The rest of the line is passed verbatim as the kernel command-line. Any modules must be reloaded after using this command.

9 They’re loaded the same way, so we will refer to the Stage 1.5 as a Stage 2 from now on.

This command also accepts the option ‘--type’ so that you can specify the kernel type of file explicitly. The argument type must be one of these: ‘netbsd’, ‘freebsd’, ‘openbsd’, ‘linux’, ‘biglinux’, and ‘multiboot’. However, you need to specify it only if you want to load a NetBSD elf kernel, because GRUB can automatically determine a kernel type in the other cases, quite safely.

The option ‘--no-mem-option’ is effective only for Linux. If the option is specified, GRUB doesn’t pass the option ‘mem=’ to the kernel.

13.3.21 lock

Command lock

Prevent normal users from executing arbitrary menu entries. You must use the command password if you really want this command to be useful (see Section 13.2.10 [password], page 26).

This command is used in a menu, as shown in this example:

     title This entry is too dangerous to be executed by normal users
     lock
     root (hd0,a)
     kernel /no-security-os

See also Chapter 9 [Security], page 17.

13.3.22 makeactive

Command makeactive

Set the active partition on the root disk to GRUB’s root device. This command is limited to primary PC partitions on a hard disk.

13.3.23 map

Command map to drive from drive Map the drive from drive to the drive to drive. This is necessary when you chain-load some operating systems, such as DOS, if such an OS resides at a non-first drive. Here is an example:

     grub> map (hd0) (hd1)
     grub> map (hd1) (hd0)

The example exchanges the order between the first hard disk and the second hard disk. See also Section 4.2.6 [DOS/Windows], page 11.

13.3.24 md5crypt

Command md5crypt

Prompt to enter a password, and encrypt it in MD5 format. The encrypted password can be used with the command password (see Section 13.2.10 [password], page 26). See also Chapter 9 [Security], page 17.

13.3.25 module

Command module file …

Load a boot module file for a Multiboot format boot image (no interpretation of the file contents are made, so that user of this command must know what the kernel in question expects). The rest of the line is passed as the module command-line, like the kernel command. You must load a Multiboot kernel image before loading any module. See also Section 13.3.26 [modulenounzip], page 35.

13.3.26 modulenounzip

Command modulenounzip file …

The same as module (see Section 13.3.25 [module], page 35), except that automatic decompression is disabled.

13.3.27 pause

Command pause message … Print the message, then wait until a key is pressed. Note that placing ^{⟨}\underline{\overline{^G}}^{⟩} (ASCII code 7) in the message will cause the speaker to emit the standard beep sound, which is useful when prompting the user to change floppies.

13.3.28 quit

Command quit

Exit from the grub shell grub (see Chapter 15 [Invoking the grub shell], page 42). This command can be used only in the grub shell.

13.3.29 reboot

Command reboot

Reboot the computer.

13.3.30 read

Command read addr

Read a 32-bit value from memory at address addr and display it in hex format.

13.3.31 root

Command root device [hdbias] Set the current root device to the device device, then attempt to mount it to get the partition size (for passing the partition descriptor in ES:ESI, used by some chainloaded boot loaders), the BSD drive-type (for booting BSD kernels using their native boot format), and correctly determine the PC partition where a BSD sub-partition is located. The optional hdbias parameter is a number to tell a BSD kernel how many BIOS drive numbers are on controllers before the current one. For example, if there is an IDE disk and a SCSI disk, and your FreeBSD root partition is on the SCSI disk, then use a ‘1’ for hdbias.

See also Section 13.3.32 [rootnoverify], page 36.

13.3.32 rootnoverify

Command rootnoverify device [hdbias] Similar to root (see Section 13.3.31 [root], page 36), but don’t attempt to mount the partition. This is useful for when an OS is outside of the area of the disk that GRUB can read, but setting the correct root device is still desired. Note that the items mentioned in root above which derived from attempting the mount will not work correctly.

13.3.33 savedefault

Command savedefault

Save the current menu entry as a default entry. Here is an example:

     default saved
     timeout 10
     title GNU/Linux
     root (hd0,0)
     kernel /boot/vmlinuz root=/dev/sda1 vga=ext
     initrd /boot/initrd
     savedefault
     title FreeBSD
     root (hd0,a)
     kernel /boot/loader
     savedefault

With this configuration, GRUB will choose the entry booted previously as the default entry. See also Section 13.1.1 [default], page 22.

13.3.34 setup

Command setup [‘--force-lba’] [‘--stage2=os_stage2_file’] [‘--prefix=dir’]

install device [image device] Set up the installation of GRUB automatically. This command uses the more flexible command install (see Section 13.3.18 [install], page 32) in the backend and installs GRUB into the device install device. If image device is specified, then find the GRUB images (see Chapter 10 [Images], page 18) in the device image device, otherwise use the current root device, which can be set by the command root. If install device is a hard disk, then embed a Stage 1.5 in the disk if possible.

The option ‘--prefix’ specifies the directory under which GRUB images are put. If it is not specified, GRUB automatically searches them in ‘/boot/grub’ and ‘/grub’.

The options ‘--force-lba’ and ‘--stage2’ are just passed to install if specified. See Section 13.3.18 [install], page 32, for more information.

13.3.35 testload

Command testload file

Read the entire contents of file in several different ways and compares them, to test the filesystem code. The output is somewhat cryptic, but if no errors are reported and the final ‘i=X, filepos=Y ’ reading has X and Y equal, then it is definitely consistent, and very likely works correctly subject to a consistent offset error. If this test succeeds, then a good next step is to try loading a kernel.

13.3.36 testvbe

Command testvbe mode

Test the VESA BIOS EXTENSION mode mode. This command will switch your video card to the graphics mode, and show an endless animation. Hit any key to return. See also Section 13.3.38 [vbeprobe], page 38.

13.3.37 uppermem

Command uppermem kbytes

Force GRUB to assume that only kbytes kilobytes of upper memory are installed. Any system address range maps are discarded.

Caution: This should be used with great caution, and should only be necessary on some old machines. GRUB’s BIOS probe can pick up all ram on all new machines the author has ever heard of. It can also be used for debugging purposes to lie to an OS.

13.3.38 vbeprobe

Command vbeprobe [mode]

Probe VESA BIOS EXTENSION information. If the mode mode is specified, show only the information about mode. Otherwise, this command lists up available VBE modes on the screen. See also Section 13.3.36 [testvbe], page 37.

14 Error messages reported by GRUB

This chapter describes error messages reported by GRUB when you encounter trouble. See Chapter 15 [Invoking the grub shell], page 42, if your problem is specific to the grub shell.

14.1 Errors reported by the Stage 1

The general way that the Stage 1 handles errors is to print an error string and then halt. Pressing ^{⟨}\underline{\overline{CTRL}}^{⟩}-^{⟨}\underline{\overline{ALT}}^{⟩}-^{⟨}\underline{\overline{DEL}}^{⟩} will reboot.

The following is a comprehensive list of error messages for the Stage 1:

Hard Disk Error The stage2 or stage1.5 is being read from a hard disk, and the attempt to determine the size and geometry of the hard disk failed.

Floppy Error The stage2 or stage1.5 is being read from a floppy disk, and the attempt to determine the size and geometry of the floppy disk failed. It’s listed as a separate error since the probe sequence is different than for hard disks.

Read Error A disk read error happened while trying to read the stage2 or stage1.5.

Geom Error The location of the stage2 or stage1.5 is not in the portion of the disk supported directly by the BIOS read calls. This could occur because the BIOS translated geometry has been changed by the user or the disk is moved to another machine or controller after installation, or GRUB was not installed using itself (if it was, the Stage 2 version of this error would have been seen during that process and it would not have completed the install).

14.2 Errors reported by the Stage 1.5

The general way that the Stage 1.5 handles errors is to print an error number in the form Error num and then halt. Pressing ^{⟨}\underline{\overline{CTRL}}^{⟩}-^{⟨}\underline{\overline{ALT}}^{⟩}-^{⟨}\underline{\overline{DEL}}^{⟩} will reboot.

The error numbers correspond to the errors reported by Stage 2. See Section 14.3 [Stage2 errors], page 39.

14.3 Errors reported by the Stage 2

The general way that the Stage 2 handles errors is to abort the operation in question, print an error string, then (if possible) either continue based on the fact that an error occurred or wait for the user to deal with the error.

The following is a comprehensive list of error messages for the Stage 2 (error numbers for the Stage 1.5 are listed before the colon in each description):

1 : Filename must be either an absolute filename or blocklist This error is returned if a file name is requested which doesn’t fit the syntax/rules listed in the Chapter 11 [Filesystem], page 19.

2 : Bad file or directory type This error is returned if a file requested is not a regular file, but something like a symbolic link, directory, or FIFO.

3 : Bad or corrupt data while decompressing file This error is returned if the run-length decompression code gets an internal error. This is usually from a corrupt file.

4 : Bad or incompatible header in compressed file This error is returned if the file header for a supposedly compressed file is bad.

5 : Partition table invalid or corrupt This error is returned if the sanity checks on the integrity of the partition table fail. This is a bad sign.

6 : Mismatched or corrupt version of stage1/stage2 This error is returned if the install command is pointed to incompatible or corrupt versions of the stage1 or stage2. It can’t detect corruption in general, but this is a sanity check on the version numbers, which should be correct.

7 : Loading below 1MB is not supported This error is returned if the lowest address in a kernel is below the 1MB boundary. The Linux zImage format is a special case and can be handled since it has a fixed loading address and maximum size.

8 : Kernel must be loaded before booting This error is returned if GRUB is told to execute the boot sequence without having a kernel to start.

9 : Unknown boot failure This error is returned if the boot attempt did not succeed for reasons which are unknown.

10 : Unsupported Multiboot features requested This error is returned when the Multiboot features word in the Multiboot header requires a feature that is not recognized. The point of this is that the kernel requires special handling which GRUB is likely unable to provide.

11 : Unrecognized device string This error is returned if a device string was expected, and the string encountered didn’t fit the syntax/rules listed in the Chapter 11 [Filesystem], page 19.

12 : Invalid device requested This error is returned if a device string is recognizable but does not fall under the other device errors.

13 : Invalid or unsupported executable format This error is returned if the kernel image being loaded is not recognized as Multiboot or one of the supported native formats (Linux zImage or bzImage, FreeBSD, or NetBSD).

14 : Filesystem compatibility error, cannot read whole file Some of the filesystem reading code in GRUB has limits on the length of the files it can read. This error is returned when the user runs into such a limit.

15 : File not found This error is returned if the specified file name cannot be found, but everything else (like the disk/partition info) is OK.

16 : Inconsistent filesystem structure This error is returned by the filesystem code to denote an internal error caused by the sanity checks of the filesystem structure on disk not matching what it expects. This is usually caused by a corrupt filesystem or bugs in the code handling it in GRUB.

17 : Cannot mount selected partition This error is returned if the partition requested exists, but the filesystem type cannot be recognized by GRUB.

18 : Selected cylinder exceeds maximum supported by BIOS This error is returned when a read is attempted at a linear block address beyond the end of the BIOS translated area. This generally happens if your disk is larger than the BIOS can handle (512MB for (E)IDE disks on older machines or larger than 8GB in general).

19 : Linux kernel must be loaded before initrd This error is returned if the initrd command is used before loading a Linux kernel. Similar to the above error, it only makes sense in that case anyway.

20 : Multiboot kernel must be loaded before modules This error is returned if the module load command is used before loading a Multiboot kernel. It only makes sense in this case anyway, as GRUB has no idea how to communicate the presence of location of such modules to a non-Multiboot-aware kernel.

21 : Selected disk does not exist This error is returned if the device part of a device- or full file name refers to a disk or BIOS device that is not present or not recognized by the BIOS in the system.

22 : No such partition This error is returned if a partition is requested in the device part of a deviceor full file name which isn’t on the selected disk.

23 : Error while parsing number This error is returned if GRUB was expecting to read a number and encountered bad data.

24 : Attempt to access block outside partition This error is returned if a linear block address is outside of the disk partition. This generally happens because of a corrupt filesystem on the disk or a bug in the code handling it in GRUB (it’s a great debugging tool).

25 : Disk read error This error is returned if there is a disk read error when trying to probe or read data from a particular disk.

26 : Too many symbolic links This error is returned if the link count is beyond the maximum (currently 5), possibly the symbolic links are looped.

27 : Unrecognized command This error is returned if an unrecognized command is entered into the commandline or in a boot sequence section of a configuration file and that entry is selected.

28 : Selected item cannot fit into memory This error is returned if a kernel, module, or raw file load command is either trying to load its data such that it won’t fit into memory or it is simply too big.

29 : Disk write error This error is returned if there is a disk write error when trying to write to a particular disk. This would generally only occur during an install of set active partition command.

30 : Invalid argument This error is returned if an argument specified to a command is invalid.

31 : File is not sector aligned This error may occur only when you access a ReiserFS partition by block-lists (e.g. the command install). In this case, you should mount the partition with the ‘-o notail’ option.

32 : Must be authenticated This error is returned if you try to run a locked entry. You should enter a correct password before running such an entry.

33 : Serial device not configured This error is returned if you try to change your terminal to a serial one before initializing any serial device.

34 : No spare sectors on the disk This error is returned if a disk doesn’t have enough spare space. This happens when you try to embed Stage 1.5 into the unused sectors after the MBR, but the first partition starts right after the MBR or they are used by EZ-BIOS.

15 Invoking the grub shell

This chapter documents the grub shell grub. Note that the grub shell is an emulator; it doesn’t run under the native environment, so it sometimes does something wrong. Therefore, you shouldn’t trust it too much. If there is anything wrong with it, don’t hesitate to try the native GRUB environment, especially when it guesses a wrong map between BIOS drives and OS devices.

15.1 Introduction into the grub shell

You can use the command grub for installing GRUB under your operating systems and for a testbed when you add a new feature into GRUB or when fix a bug. grub is almost the same as the Stage 2, and, in fact, it shares the source code with the Stage 2 and you can use the same commands (see Chapter 13 [Commands], page 22) in grub. It is emulated by replacing BIOS calls with UNIX system calls and libc functions.

The command grub accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘--verbose’

Print some verbose messages for debugging purpose.

‘--device-map=file’

Use the device map file file. The format is described in Section 15.3 [Device map], page 43.

‘--no-floppy’

Do not probe any floppy drive. This option has no effect if the option ‘--device-map’ is specified (see Section 15.3 [Device map], page 43).

‘--probe-second-floppy’

Probe the second floppy drive. If this option is not specified, the grub shell does not probe it, as that sometimes takes a long time. If you specify the device map file (see Section 15.3 [Device map], page 43), the grub shell just ignores this option.

‘--config-file=file’

Read the configuration file file instead of ‘/boot/grub/menu.lst’. The format is the same as the normal GRUB syntax. See Chapter 11 [Filesystem], page 19, for more information.

‘--boot-drive=drive’

Set the stage2 boot drive to drive. This argument should be an integer (decimal, octal or hexadecimal).

‘--install-partition=par’

Set the stage2 install partition to par. This argument should be an integer (decimal, octal or hexadecimal).

‘--no-config-file’

Do not use the configuration file even if it can be read.

‘--no-curses’

Do not use the curses interface even if it is available.

‘--batch’ This option has the same meaning as ‘--no-config-file --no-curses’.
‘--read-only’

Disable writing to any disk.

‘--hold’ Wait until a debugger will attach. This option is useful when you want to debug the startup code.

15.2 How to install GRUB via grub

The installation procedure is the same as under the native Stage 2. See Chapter 3 [Installation], page 5, for more information. The command grub-specific information is described here.

What you should be careful about is buffer cache. grub makes use of raw devices instead of filesystems that your operating systems serve, so there exists a potential problem that some cache inconsistency may corrupt your filesystems. What we recommend is:

• If you can unmount drives to which GRUB may write any amount of data, unmount them before running grub.

• If a drive cannot be unmounted but can be mounted with the read-only flag, mount it in read-only mode. That should be secure.

• If a drive must be mounted with the read-write flag, make sure that any activity is not being done on it during running the command grub.

• Reboot your operating system as soon as possible. Probably that is not required if you follow these rules above, but reboot is the most secure way.

In addition, enter the command quit when you finish the installation. That is very important because quit makes the buffer cache consistent. Do not push ^{⟨}\underline{\overline{C-c}}^{⟩}.

If you want to install GRUB non-interactively, specify ‘--batch’ option in the command-line. This is a simple example:

     #!/bin/sh
     # Use /usr/sbin/grub if you are on an older system.
     /sbin/grub --batch <<EOT 1>/dev/null 2>/dev/null
     root (hd0,0)
     setup (hd0)
     quit
     EOT

15.3 The map between BIOS drives and OS devices

When you specify the option ‘--device-map’ (see Section 15.1 [Basic usage], page 42), the grub shell creates the device map file automatically unless it already exists. The file name ‘/boot/grub/device.map’ is preferred.

If the device map file exists, the grub shell reads it to map BIOS drives to OS devices. This file consists of lines like this:

device file

device is a drive, which syntax is the same as the one in GRUB (see Section 11.1 [Device syntax], page 19), and file is an OS’s file, which is normally a device file.

The reason why the grub shell gives you the device map file is that it cannot guess the map between BIOS drives and OS devices correctly in some environments. For example, if you exchange the boot sequence between IDE and SCSI in your BIOS, it mistakes the order.

Thus, edit the file if the grub shell makes a mistake. You can put any comments in the file if needed, as the grub shell assumes that a line is just a comment if the first character is ‘#’.

16 Invoking grub-install

The program grub-install installs GRUB on your drive by the grub shell (see Chapter 15 [Invoking the grub shell], page 42). You must specify the device name on which you want to install GRUB, like this:

grub-install install device The device name install device is an OS device name or a GRUB device name. grub-install accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘--force-lba’

Force GRUB to use LBA mode even for a buggy BIOS. Use this option only if your BIOS doesn’t work in LBA mode even though it supports LBA mode.

‘--root-directory=dir’

Install GRUB images under the directory dir instead of the root directory. This option is useful when you want to install GRUB into a separate partition or a removable disk. Here is an example when you have a separate boot partition which is mounted on ‘/boot’:

               grub-install --root-directory=/boot '(hd0)'
‘--grub-shell=file’

Use file as the grub shell. You can append arbitrary options to file after the file name, like this:

               grub-install --grub-shell="grub --read-only" /dev/fd0
‘--recheck’

Recheck the device map, even if ‘/boot/grub/device.map’ already exists. You should use this option whenever you add/remove a disk into/from your computer.

17 Invoking grub-md5-crypt

The program grub-md5-crypt encrypts a password in MD5 format. This is just a frontend of the grub shell (see Chapter 15 [Invoking the grub shell], page 42). Passwords encrypted by this program can be used with the command password (see Section 13.2.10 [password], page 26).

grub-md5-crypt accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version information and exit.

‘--grub-shell=file’

Use file as the grub shell.

18 Invoking mbchk

The program mbchk checks for the format of a Multiboot kernel. We recommend using this program before booting your own kernel by GRUB.

mbchk accepts the following options:

‘--help’ Print a summary of the command-line options and exit.

‘--version’

Print the version number of GRUB and exit.

‘--quiet’ Suppress all normal output.

Appendix A Frequently asked questions

How does GNU GRUB differ from Erich’s original GRUB?

GNU GRUB is the successor of Erich’s great GRUB. He couldn’t work on GRUB because of some other tasks, so the current maintainers Yoshinori K. Okuji and Gordon Matzigkeit took over the maintainership, and opened the development in order for everybody to participate it. Technically speaking, GNU GRUB has many features that are not seen in the original GRUB. For example, GNU GRUB can be installed on UNIX-like operating system (such as GNU/Hurd) via the grub shell ‘/sbin/grub’ (or ‘/usr/sbin/grub’ on older systems), it supports Logical Block Address (LBA) mode that solves the 1024 cylinders problem, and ^{⟨}\underline{\overline{TAB}}^{⟩} completes a file name when it’s unique. Of course, many bug fixes are done as well, so it is recommended to use GNU GRUB.

Can GRUB boot my operating system from over 8GB hard disks?

That depends on your BIOS and your operating system. You must make sure that your drive is accessible in LBA mode. Generally, that is configurable in BIOS setting utility. Read the manual for your BIOS for more information.

Furthermore, some operating systems (i.e. DOS) cannot access any large disk, so the problem is not solved by any kind of boot loader. GNU/Hurd and GNU/Linux can surely boot from such a large disk.

Can I put Stage2 into a partition which is over 1024 cylinders?

Yes, if your BIOS supports the LBA mode.

How to create a GRUB boot floppy with the menu interface?

The easiest way is:

1. Create filesystem in your floppy disk. For example:

         $ mke2fs /dev/fd0

2. Mount it on somewhere, say, ‘/mnt’.

3. Copy the GRUB images to ‘/mnt/boot/grub’. Only ‘stage1’, ‘stage2’ and ‘menu.lst’ are necessary. You may not copy stage1.5s.

4. Run the following command (substitute ‘/usr/sbin/grub’ for

‘/sbin/grub’ if you are using an older system):

         $ /sbin/grub --batch <<EOT
         root (fd0)
         setup (fd0)
         quit
         EOT

How to specify a partition?

See Section 11.1 [Device syntax], page 19.

GRUB does not recognize my GNU/Hurd partition.

I don’t know why, but the authors of FDISK programs have assigned the partition type ‘0x63’ to the GNU Hurd incorrectly. So use ‘0x83’ if the partition contains ext2fs filesystem, and use ‘0xA5’ if the partition contains ffs filesystem, whether the partition owner is the Hurd or not. We will use ‘0x63’ for the GNU Hurd filesystem that has not been implemented yet.

I’ve installed a recent version of binutils, but GRUB still crashes.

Please check for the version of your binutils by this command:

     ld -v

This will show two versions, but only the latter is important. If the version is identical with what you have installed, the installation was not bad. Well, please try:

     gcc -Wl,-v 2>&1 | grep "GNU ld"

If this is not identical with the result above, you should specify the directory where you have installed binutils for the script configure, like this:

     ./configure --with-binutils=/usr/local/bin

If you follow the instructions above but GRUB still crashes, probably there is a serious bug in GRUB. See Appendix C [Reporting bugs], page 49.

GRUB hangs up when accessing my SCSI disk.

Check if you have turned on the support for INT 13 extension (LBA). If so, disable the support and see if GRUB can now access your SCSI disk. This will make it clear that your SCSI BIOS sucks.

For now, we know the following doesn’t provide working LBA mode:

Adaptec AIC-7880

In the case where you have such a SCSI controller unfortunately, you cannot use the LBA mode, though GRUB still works fine in the CHS mode (so the well-known 1024 cylinders problem comes again to you). Caution: Actually it has not been verified yet if this bug is due to the SCSI BIOS or GRUB itself, frankly speaking. Because the developers haven’t seen it by their own eyes. This is why it is desirable that you investigate the cause seriously if you have the skill.

How can I specify an arbitrary memory size to Linux?

Pass a ‘mem=’ option to your Linux kernel, like this:

     grub> kernel /vmlinuz mem=128M

You may pass other options in the same way. See See Section 4.2.2 [GNU/Linux], page 9, for more details.

I have a separate boot partition and GRUB doesn’t recognize it.

This is often reported as a bug, but this is not a bug really. This is a feature. Because GRUB is a boot loader and it normally runs under no operating system, it doesn’t know where a partition is mounted under your operating systems. So, if you have the partition ‘/boot’ and you install GRUB images into the directory ‘/boot/grub’, GRUB recognizes that the images lies under the directory ‘/grub’ but not ‘/boot/grub’. That’s fine, since there is no guarantee that all of your operating systems mount the same partition as ‘/boot’. There are several solutions for this situation.

1. Install GRUB into the directory ‘/boot/boot/grub’ instead of ‘/boot/grub’. This may sound ugly but should work fine.

2. Create a symbolic link before installing GRUB, like ‘cd /boot && ln -s . boot’. This works only if the filesystem of the boot partition supports symbolic links and GRUB supports the feature as well.

3. Install GRUB with the command install, to specify the paths of GRUB

images explicitly. Here is an example:

         grub> root (hd0,1)
         grub> install /grub/stage1 d (hd0) /grub/stage2 p /grub/menu.lst

How to uninstall GRUB from my hard disk drive?

There is no concept uninstall in boot loaders, because if you uninstall a boot loader, an unbootable machine would simply remain. So all you need to do is overwrite another boot loader you like to your disk, that is, install the boot loader without uninstalling GRUB. For example, if you want to install the boot loader for Windows, just run FDISK /MBR on Windows. If you want to install LILO^{10}, run /sbin/lilo on GNU/Linux.

10 I can’t imagine why you want to do such a thing, though

GRUB hangs when accessing my large IDE disk.

If your disk is bigger than 32GB, probably updating your mainboard BIOS will solve your problem. This bug is well-known and most vendors should provide fixed versions. For example, if you have ASUS-P3BF, upgrading the BIOS to V1007beta1 or later can fix it. Please ask your vendor, for more information.

Why don’t Linux, FreeBSD, NetBSD, etc. become Multiboot-compliant?

Please ask the relevant maintainers. If all free kernels were Multiboot-compliant (see section “Motivation” in The Multiboot Specification), the world would be an utopia…

Appendix B How to obtain and build GRUB

Caution: GRUB requires binutils-2.9.1.0.23 or later because the GNU assembler has been changed so that it can produce real 16bits machine code between

2.9.1 and 2.9.1.0.x. See http://sources.redhat.com/binutils/, to obtain

information on how to get the latest version.

GRUB is available from the GNU alpha archive site ftp://alpha.gnu.org/gnu/grub or any of its mirrors. The file will be named grub-version.tar.gz. The current version is 0.92, so the file you should grab is:

 ftp://alpha.gnu.org/gnu/grub/grub-0.92.tar.gz

To unbundle GRUB use the instruction:

zcat grub-0.92.tar.gz | tar xvf -

which will create a directory called ‘grub-0.92’ with all the sources. You can look at the file ‘INSTALL’ for detailed instructions on how to build and install GRUB, but you should be able to just do:

cd grub-0.92
./configure
make install

This will install the grub shell ‘grub’ (see Chapter 15 [Invoking the grub shell], page 42), the Multiboot checker ‘mbchk’ (see Chapter 18 [Invoking mbchk], page 45), and the GRUB images. This will also install the GRUB manual.

Also, the latest version is available from the CVS. The repository is:

 :pserver:anoncvs@subversions.gnu.org:/cvsroot/grub

and the module is:

 grub

The password for anoncvs is empty. So the instruction is:

cvs -d :pserver:anoncvs@subversions.gnu.org:/cvsroot/grub login
Password: ^{⟨}\underline{\overline{ENTER}}^{⟩}
cvs -d :pserver:anoncvs@subversions.gnu.org:/cvsroot/grub co grub

Appendix C Reporting bugs

This is the guideline of how to report bugs. Take a look at this list below before you

send e-mail to bug-grub@gnu.org:

1. Before unsettled, read this manual through and through, in particular See Appendix A [FAQ], page 45.

2. Always mention the information on your GRUB. The version number and the configuration are quite important. If you build it yourself, write the options specified to the configure script and your operating system, including the versions of gcc and binutils.

3. If you get troubled with the installation, inform us of how you installed GRUB. Don’t omit error messages, if any. Just ‘GRUB hangs up when it boots’ is not enough. The information on your hardware is also essential. These are especially important: the geometries and the partition tables of your hard disk drives and your BIOS.

4. If GRUB cannot boot your operating system, write down all on the screen. Don’t

paraphrase them, like ‘The foo OS crashes with GRUB, even though it can boot

with the bar boot loader fine’. Mention the commands you executed, the messages printed by them, and information on your operating system including the version number.

5. Explain what you wanted to do. It is very useful to know your purpose and your wish, and how GRUB didn’t satisfy you.

6. If you can investigate the problem yourself, please do. That will give you and us much more information on the problem. Attaching a patch is even better. When you attach a patch, make the patch in unified diff format, and write ChangeLog entries. But, even when you make a patch, don’t forget to explain the problem, so that we can understand what your patch is for.

7. Write down anything that you think might be related. Please understand that we often need to reproduce the same problem you encounterred in our environment. So your information should be sufficient for us to do the same thing—Don’t forget that we cannot see your computer directly. If you are not sure whether to state a fact or leave it out, state it! Reporting too many things is quite better than omitting an important thing.

If you realize the guideline above, send e-mail to bug-grub@gnu.org, and we will try to fix the bugs.

19 Where GRUB will go

Here are some ideas that might happen in the future:

• Support dynamic loading.

• Add real memory management.

• Add a real scripting language.

• Support internationalization.

• Support other architectures than i386-pc.

See the file ‘TODO’ in the source distribution, for more information.

20 Hacking GRUB

This chapter documents the user-invisible aspect of GRUB.

As a general rule of software development, it is impossible to keep the descriptions of the internals up-to-date, and it is quite hard to document everything. So refer to the source code, whenever you are not satisfied with this documentation. Please assume that this gives just hints to you.

20.1 The memory map of various components

GRUB consists of two distinct components, called stages, which are loaded at different times in the boot process. Because they run mutual-exclusively, sometimes a memory area overlaps with another memory area. And, even in one stage, a single memory area can be used for various purposes, because their usages are mutually exclusive.

Here is the memory map of the various components:

0 to 4K-1 BIOS and real mode interrupts

0x07BE to 0x07FF Partition table passed to another boot loader

down from 8K-1 Real mode stack

0x2000 to ?

The optional Stage 1.5 is loaded here

0x2000 to 0x7FFF Command-line buffer for Multiboot kernels and modules

0x7C00 to 0x7DFF Stage 1 is loaded here by BIOS or another boot loader

0x7F00 to 0x7F42 LBA drive parameters

0x8000 to ?

Stage2 is loaded here

The end of Stage 2 to 416K-1 Heap, in particular used for the menu

down from 416K-1 Protected mode stack

416K to 448K-1 Filesystem buffer

448K to 479.5K-1 Raw device buffer

479.5K to 480K-1 512-byte scratch area 480K to 512K-1

Buffers for various functions, such as password, command-line, cut and paste, and completion.

The last 1K of lower memory

Disk swapping code and data See the file ‘stage2/shared.h’, for more information.

20.2 Embedded variables in GRUB

Stage 1 and Stage 2 have embedded variables whose locations are well-defined, so that the installation can patch the binary file directly without recompilation of the stages.

In Stage 1, these are defined: 0x3E The version number (not GRUB’s, but the installation mechanism’s). 0x40 The boot drive. If it is 0xFF, use a drive passed by BIOS. 0x41 The flag for if forcing LBA. 0x42 The starting address of Stage 2. 0x44 The first sector of Stage 2. 0x48 The starting segment of Stage 2. 0x1FE The signature (0xAA55).

See the file ‘stage1/stage1.S’, for more information.

In the first sector of Stage 1.5 and Stage 2, the block lists are recorded between firstlist and lastlist. The address of lastlist is determined when assembling the file

‘stage2/start.S’.

The trick here is that it is actually read backward, and the first 8-byte block list is not read here, but after the pointer is decremented 8 bytes, then after reading it, it decrements again, reads, and so on, until it is finished. The terminating condition is when the number of sectors to be read in the next block list is zero.

The format of a block list can be seen from the example in the code just before the firstlist label. Note that it is always from the beginning of the disk, but not relative to the partition boundaries.

In the second sector of Stage 1.5 and Stage 2, these are defined: 0x6 The version number (likewise, the installation mechanism’s). 0x8 The installed partition. 0xC The saved entry number. 0x10 The identifier. 0x11 The flag for if forcing LBA. 0x12 The version string (GRUB’s). 0x12 + the length of the version string

The name of a configuration file. See the file ‘stage2/asm.S’, for more information.

20.3 The generic interface for filesystems

For any particular partition, it is presumed that only one of the normal filesystems such as FAT, FFS, or ext2fs can be used, so there is a switch table managed by the functions in ‘disk_io.c’. The notation is that you can only mount one at a time.

The block list filesystem has a special place in the system. In addition to the normal filesystem (or even without one mounted), you can access disk blocks directly (in the indicated partition) via the block list notation. Using the block list filesystem doesn’t effect any other filesystem mounts.

The variables which can be read by the filesystem backend are:

current_drive

The current BIOS drive number (numbered from 0, if a floppy, and numbered from 0x80, if a hard disk).

current_partition

The current partition number.

current_slice

The current partition type.

saved_drive

The drive part of the root device.

saved_partition

The partition part of the root device.

part_start

The current partition starting address, in sectors.

part_length

The current partition length, in sectors.

print_possibilities

True when the dir function should print the possible completions of a file, and false when it should try to actually open a file of that name.

FSYS_BUF Filesystem buffer which is 32K in size, to use in any way which the filesystem backend desires.

The variables which need to be written by a filesystem backend are:

filepos The current position in the file, in sectors.

Caution: the value of filepos can be changed out from under the filesystem code in the current implementation. Don’t depend on it being the same for later calls into the backend code!

filemax The length of the file.

disk_read_func

The value of disk read hook only during reading of data for the file, not any other fs data, inodes, FAT tables, whatever, then set to NULL at all other times (it will be NULL by default). If this isn’t done correctly, then the testload and install commands won’t work correctly.

The functions expected to be used by the filesystem backend are:

devread Only read sectors from within a partition. Sector 0 is the first sector in the partition.

grub_read

If the backend uses the block list code, then grub_read can be used, after setting block file to 1.

print_a_completion

If print possibilities is true, call print_a_completion for each possible file name. Otherwise, the file name completion won’t work.

The functions expected to be defined by the filesystem backend are described at least moderately in the file ‘filesys.h’. Their usage is fairly evident from their use in the functions in ‘disk_io.c’, look for the use of the fsys table array.

Caution: The semantics are such that then ‘mount’ing the filesystem, presume the filesystem buffer FSYS_BUF is corrupted, and (re-)load all important contents. When opening and reading a file, presume that the data from the ‘mount’ is available, and doesn’t get corrupted by the open/read (i.e. multiple opens and/or reads will be done with only one mount if in the same filesystem).

20.4 The generic interface for built-ins

GRUB built-in commands are defined in a uniformal interface, whether they are menu-specific or can be used anywhere. The definition of a builtin command consists of two parts: the code itself and the table of the information.

The code must be a function which takes two arguments, a command-line string and flags, and returns an ‘int’ value. The flags argument specifies how the function is called, using a bit mask. The return value must be zero if successful, otherwise non-zero. So it is normally enough to return errnum.

The table of the information is represented by the structure struct builtin, which contains the name of the command, a pointer to the function, flags, a short description of the command and a long description of the command. Since the descriptions are used only for help messages interactively, you don’t have to define them, if the command may not be called interactively (such as title).

The table is finally registered in the table builtin table, so that run_script and enter_cmdline can find the command. See the files ‘cmdline.c’ and ‘builtins.c’, for more details.

20.5 The bootstrap mechanism used in GRUB

The disk space can be used in a boot loader is very restricted because a MBR (see Section 20.9 [MBR], page 55) is only 512 bytes but it also contains a partition table (see Section 20.10 [Partition table], page 55) and a BPB. So the question is how to make a boot loader code enough small to be fit in a MBR.

However, GRUB is a very large program, so we break GRUB into 2 (or 3) distinct components, Stage 1 and Stage 2 (and optionally Stage 1.5). See Section 20.1 [Memory map], page 50, for more information.

We embed Stage 1 in a MBR or in the boot sector of a partition, and place Stage 2 in a filesystem. The optional Stage 1.5 can be installed in a filesystem, in the boot loader area in a FFS or a ReiserFS, and in the sectors right after a MBR, because Stage 1.5 is enough small and the sectors right after a MBR is normally an unused region. The size of this region is the number of sectors per head minus 1.

Thus, all Stage1 must do is just load Stage2 or Stage1.5. But even if Stage 1 needs not to support the user interface or the filesystem interface, it is impossible to make Stage 1 less than 400 bytes, because GRUB should support both the CHS mode and the LBA mode (see Section 20.8 [Low-level disk I/O], page 54).

The solution used by GRUB is that Stage 1 loads only the first sector of Stage 2 (or Stage 1.5) and Stage 2 itself loads the rest. The flow of Stage 1 is:

1. Initialize the system briefly.

2. Detect the geometry and the accessing mode of the loading drive.

3. Load the first sector of Stage 2.

4. Jump to the starting address of the Stage 2.

The flow of Stage 2 (and Stage 1.5) is:

1. Load the rest of itself to the real starting address, that is, the starting address plus 512 bytes. The block lists are stored in the last part of the first sector.

2. Long jump to the real starting address.

Note that Stage 2 (or Stage 1.5) does not probe the geometry or the accessing mode of the loading drive, since Stage 1 has already probed them.

20.6 How to probe I/O ports used by INT 13H

FIXME: I will write this chapter after implementing the new technique.

20.7 How to detect all installed RAM

FIXME: I doubt if Erich didn’t write this chapter only himself wholly, so I will rewrite this chapter.

20.8 INT 13H disk I/O interrupts

FIXME: I’m not sure where some part of the original chapter is derived, so I will rewrite this chapter.

20.9 The structure of Master Boot Record

FIXME: Likewise.

20.10 The format of partition tables

FIXME: Probably the original chapter is derived from "How It Works", so I will rewrite this chapter.

20.11 Where and how you should send patches

When you write patches for GRUB, please send them to the mailing list bug-grub@gnu.org. Here is the list of items of which you should take care:

• Please make your patch as small as possible. Generally, it is not a good thing to make one big patch which changes many things. Instead, segregate features and produce many patches.

• Use as late code as possible, for the original code. The CVS repository always has the current version (see Appendix B [Obtaining and Building GRUB], page 48).

• Write ChangeLog entries. See section “Change Logs” in GNU Coding Standards, if you don’t know how to write ChangeLog.

• Make patches in unified diff format. ‘diff -urN’ is appropriate in most cases.

• Don’t make patches reversely. Reverse patches are difficult to read and use.

• Be careful enough of the license term and the copyright. Because GRUB is under GNU General Public License, you may not steal code from software whose license is incompatible against GPL. And, if you copy code written by others, you must not ignore their copyrights. Feel free to ask GRUB maintainers, whenever you are not sure what you should do.

• If your patch is too large to send in e-mail, put it at somewhere we can see. Usually, you shouldn’t send e-mail over 20K.

Index

B

blocklist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 boot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 bootp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

C

cat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 chainloader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 cmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 configfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 current_drive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 current_partition . . . . . . . . . . . . . . . . . . . . . . . . . . 52 current_slice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

D

debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 default. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 devread. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 dhcp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 disk_read_func . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 displayapm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 displaymem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

56 GRUB Manual

E

embed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

F

fallback. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 filemax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 filepos. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 fstest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 FSYS_BUF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

G

geometry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 grub_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

H

halt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 hiddenmenu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 hide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

I

ifconfig. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 impsprobe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 initrd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 install. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 ioprobe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

K

kernel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

L

lock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

M

makeactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 md5crypt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 modulenounzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

P

pager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 part_length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 part_start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 partnew. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 parttype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 password. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 pause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 print_a_completion . . . . . . . . . . . . . . . . . . . . . . . . . 53 print_possibilities. . . . . . . . . . . . . . . . . . . . . . . . 52

Q

quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

R

rarp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 reboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 root . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 rootnoverify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

S

saved_drive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 saved_partition . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 savedefault . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 serial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 setkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

T

terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 testload. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 testvbe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 tftpserver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 timeout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

U

unhide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 uppermem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

V

vbeprobe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

Table of Contents

1 Introduction to GRUB . . . . . . . . . . . . . . . . . . . . . 1

1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 History of GRUB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 GRUB features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 The role of a boot loader. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Naming convention . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Creating a GRUB boot floppy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Installing GRUB natively . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 Installing GRUB using grub-install . . . . . . . . . . . . . . . . . . . . . . . 7

4 Booting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.1 How to boot operating systems . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.1.1 How to boot an OS directly with GRUB. . . . . . . . . . 8 4.1.2 Load another boot loader to boot unsupported

operating systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4.2 Some caveats on OS-specific issues . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2.1 GNU/Hurd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2.2 GNU/Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2.3 FreeBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.2.4 NetBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.2.5 OpenBSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.2.6 DOS/Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2.7 SCO UnixWare . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5 Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6 Downloading OS images from a network … 14

6.1 How to set up your network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 6.2 Booting from a network. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

7 Using GRUB via a serial line . . . . . . . . . . . . . . 15

8 Embedding a configuration file into GRUB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

9 Protecting your computer from cracking … 17 10 GRUB image files . . . . . . . . . . . . . . . . . . . . . . . 18

11 Filesystem syntax and semantics. . . . . . . . . . 19

11.1 How to specify devices. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 11.2 How to specify files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 11.3 How to specify block lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

12 GRUB’s user interface . . . . . . . . . . . . . . . . . . . 20

12.1 The flexible command-line interface. . . . . . . . . . . . . . . . . . . . . 20 12.2 The simple menu interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 12.3 Editing a menu entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 12.4 The hidden menu interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

13 The list of available commands . . . . . . . . . . . 22

13.1 The list of commands for the menu only . . . . . . . . . . . . . . . . 22 13.1.1 default. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 13.1.2 fallback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 13.1.3 hiddenmenu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 13.1.4 timeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 13.1.5 title . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

13.2 The list of general commands . . . . . . . . . . . . . . . . . . . . . . . . . . 23 13.2.1 bootp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 13.2.2 color. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 13.2.3 device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 13.2.4 dhcp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 13.2.5 hide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 13.2.6 ifconfig . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 13.2.7 pager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 13.2.8 partnew . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 13.2.9 parttype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.2.10 password. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.2.11 rarp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.2.12 serial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.2.13 setkey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 13.2.14 terminal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 13.2.15 tftpserver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 13.2.16 unhide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

13.3 The list of command-line and menu entry commands. . . . . 29 13.3.1 blocklist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 13.3.2 boot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 13.3.3 cat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 13.3.4 chainloader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 13.3.5 cmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 13.3.6 configfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 13.3.7 debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 13.3.8 displayapm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 13.3.9 displaymem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 13.3.10 embed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 13.3.11 find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 13.3.12 fstest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 13.3.13 geometry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 13.3.14 halt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 13.3.15 help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 13.3.16 impsprobe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 13.3.17 initrd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 13.3.18 install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 13.3.19 ioprobe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 13.3.20 kernel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 13.3.21 lock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 13.3.22 makeactive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 13.3.23 map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 13.3.24 md5crypt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 13.3.25 module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 13.3.26 modulenounzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 13.3.27 pause . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 13.3.28 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 13.3.29 reboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 13.3.30 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 13.3.31 root . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 13.3.32 rootnoverify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 13.3.33 savedefault . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 13.3.34 setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 13.3.35 testload. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 13.3.36 testvbe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 13.3.37 uppermem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 13.3.38 vbeprobe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

14 Error messages reported by GRUB . . . . . . . 38

14.1 Errors reported by the Stage 1 . . . . . . . . . . . . . . . . . . . . . . . . . 38 14.2 Errors reported by the Stage 1.5. . . . . . . . . . . . . . . . . . . . . . . . 38 14.3 Errors reported by the Stage 2 . . . . . . . . . . . . . . . . . . . . . . . . . 39

15 Invoking the grub shell. . . . . . . . . . . . . . . . . . . 42

15.1 Introduction into the grub shell . . . . . . . . . . . . . . . . . . . . . . . . 42 15.2 How to install GRUB via grub . . . . . . . . . . . . . . . . . . . . . . . . . 43 15.3 The map between BIOS drives and OS devices . . . . . . . . . . 43

16 Invoking grub-install . . . . . . . . . . . . . . . . . . . . . 44

17 Invoking grub-md5-crypt . . . . . . . . . . . . . . . . . 45

18 Invoking mbchk . . . . . . . . . . . . . . . . . . . . . . . . . 45 Appendix A Frequently asked questions. . . . . . 45

Appendix B How to obtain and build GRUB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

Appendix C Reporting bugs. . . . . . . . . . . . . . . . . 49

19 Where GRUB will go . . . . . . . . . . . . . . . . . . . . 49

20 Hacking GRUB. . . . . . . . . . . . . . . . . . . . . . . . . . 50

20.1 The memory map of various components . . . . . . . . . . . . . . . . 50 20.2 Embedded variables in GRUB. . . . . . . . . . . . . . . . . . . . . . . . . . 51 20.3 The generic interface for filesystems . . . . . . . . . . . . . . . . . . . . 52 20.4 The generic interface for built-ins. . . . . . . . . . . . . . . . . . . . . . . 53 20.5 The bootstrap mechanism used in GRUB . . . . . . . . . . . . . . . 54 20.6 How to probe I/O ports used by INT 13H. . . . . . . . . . . . . . . 54 20.7 How to detect all installed RAM . . . . . . . . . . . . . . . . . . . . . . . 54 20.8 INT 13H disk I/O interrupts . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 20.9 The structure of Master Boot Record . . . . . . . . . . . . . . . . . . . 55 20.10 The format of partition tables. . . . . . . . . . . . . . . . . . . . . . . . . 55 20.11 Where and how you should send patches . . . . . . . . . . . . . . . 55

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55